### v1

# Learning Diffusion Policies for Robotic Manipulation of Timber Joinery under Fabrication Uncertainty

Abstract
Construction uncertainties such as fabrication inaccuracies and material imperfections pose a significant challenge to contact-rich robotic manipulation by hindering precise and robust assembly. In this paper, we explore the performance and robustness of diffusion policy learning as a promising solution for contact-sensitive robotic assembly at construction scale, using timber mortise and tenon joints as a case study. A two-phase study is conducted: first, to evaluate policy performance and applicability; second, to assess robustness in handling fabrication uncertainties simulated as randomized perturbations to the mortise position. The best-performing policy achieved a total average success rate of 75% with perturbations up to 10 mm, including 100% success in unperturbed cases. The results demonstrate the potential of sensory–motor diffusion policies to generalize to a wide range of complex, contact-rich assembly tasks across construction and manufacturing, advancing robotic construction under uncertainty and contributing to safer, more efficient building practices.
keywords:
Contact-rich manipulation, Robotic assembly, Fabrication uncertainty, Diffusion policy, Timber joinery, Robot learning, Construction robotics
1 Introduction
Construction plays a critical role in the global economy but continues to face long-standing challenges, including stagnant productivity growth, shortages of skilled labor, and persistent health and safety concerns Delgado et al. [2019], Wei et al. [2023]. The industry often requires workers to perform repetitive and physically demanding tasks, such as lifting and positioning heavy components, which can contribute to chronic injuries and reduced long-term workforce capacity Laukkanen [1999], Arndt et al. [2005]. Automation, and in particular robotic assembly, offers the potential to address these issues by enhancing precision, increasing efficiency, enabling mass customization, and reducing the physical burden on workers Bock [2015], Musarat et al. [2024], Chen and Adel [2025]. Recent advances in construction robotics have increasingly harnessed the precision and programmability of industrial robotic arms to execute complex and customized assemblies that would be difficult to achieve through manual methods Graser et al. [2021], Chen and Adel [2025]. Despite these advancements, translating the capabilities of industrial robots from manufacturing to the uncertain and unstructured conditions of construction, characterized by fabrication inaccuracies, material imperfections, as well as dynamic and evolving environments, remains a significant challenge, making it difficult to achieve the same levels of repeatability, quality, and precision that are standard in factory settings Adel et al. [2024], Delgado et al. [2019], Musarat et al. [2024].
The momentum in construction robotics research is especially evident in recent work on robotic timber assembly Willmann et al. [2017], Apolinarska [2018], Adel et al. [2018], Adel [2020, 2022], Chai et al. [2022], Lauer et al. [2023], where many workflows rely on planar face-to-face or butt joints, whose geometric simplicity allows the use of pre-programmed trajectories for successful assembly. In contrast, the manipulation of traditional timber joinery such as mortise and tenon, lap, or scarf joints requires contact-sensitive insertion and interlocking, demanding high precision and control of contact forces Apolinarska [2018], Leung et al. [2021]. Similar to many contact-rich robotic manipulation tasks Suomalainen et al. [2022], the assembly of these joints can lead to misalignment and jamming if not coupled with real-time control for robust error recovery Apolinarska et al. [2021], Kramberger et al. [2022]. As a result, these contact-sensitive tasks are often delegated to skilled human workers, who intervene to clamp, hammer, and adjust fits, compensating for fabrication inaccuracies and material imperfections Yang et al. [2024]. While effective in the short term, such reliance on manual intervention can break the continuity of digital production workflows, introducing delays and reducing overall efficiency. Furthermore, it increases reliance on a shrinking pool of skilled labor and entails physically demanding, labor-intensive work that can contribute to worker fatigue and injury.
Figure 1: Overview of our method. A CNN-based diffusion policy is trained conditional on end-effector pose and F/T data observations to predict sequences of robot actions.
Timber joinery is a centuries-old construction technique of geometrically interlocking timber components to create durable structural connections with minimal reliance on fasteners or adhesives Benson [1981]. These joints are valued for their enhanced structural performance, rotational stiffness, and resilience under seismic and fire conditions Fang and Mueller [2023], Fang [2021]. While they offer significant structural advantages, their complex geometry and contact-rich assembly make them particularly sensitive to the uncertainties inherent to construction tasks, including fabrication inaccuracies from cut length and angle deviations and component shifting during picking, placing, and fastening, as well as material imperfections from dimensional variations and hygrothermal expansion Adel et al. [2024]. Collectively, these factors further increase the complexity of automating the assembly of such joints in construction settings. Despite the significance of this challenge, automated assembly of traditional joinery remains nascent in robotic workflows within the Architecture, Engineering, and Construction (AEC) domain, particularly when accounting for fabrication and material uncertainty. Overcoming these limitations requires feedback-driven and closed-loop planning and control methods capable of adapting to unpredictable conditions inherent in construction workflows.
Learning-based methods have increasingly replaced traditional control-based methods for planning robot actions. Sensory–motor policy learning methods integrate sensor observations, planning, and control into a single end-to-end system, originally implemented using deep convolutional neural networks (CNNs) Levine et al. [2016]. These methods enable the robot to learn how to act in response to what it senses without requiring separately engineered modules for perception, planning, state estimation, or low-level control. These learned control policies map raw observation inputs, such as force feedback and images captured from a camera, directly to motor commands or robot actions using various learning-based methods Finn et al. [2017], Levine et al. [2018], Kalashnikov et al. [2018], Brohan et al. [2023]. Among these methods, diffusion policy learning via behavior cloning Chi et al. [2024a] has shown outstanding performance for various manipulation scenarios Chi et al. [2024b], Hou et al. [2025], Yang et al. [2025], Zhu et al. [2025]. The key advantages of diffusion policies lie in their ability to learn multi-modal action distributions, capturing multiple plausible ways to perform a task, while maintaining training stability and predicting coherent sequences of future actions Chi et al. [2024a].
The research presented in this paper tackles the challenge of contact-rich robotic manipulation under construction uncertainties, where fabrication inaccuracies and material imperfections hinder precise and robust assembly. We address this by implementing diffusion policy learning, trained on robot pose and force/torque (F/T) data from teleoperated demonstrations, to plan and control the construction-scale assembly of mortise and tenon joints as a case study (Fig. 1).
1.1 Related work
In construction with discrete elements, such as brickwork and timber or steel framing, uncertainty remains a fundamental challenge to achieving precise and robust robotic assembly. In timber framing, which serves as the case study in this research, these uncertainties arise from material imperfections (e.g., dimensional deviations, deformation, shrinkage/expansion due to moisture) and fabrication inaccuracies (e.g., cut length and angle errors, pose deviations from calibration or handling). If unaccounted for, these variabilities accumulate during assembly, leading to misalignments, part collisions, and robot failures when trajectories are executed directly from as-planned digital models  Adel et al. [2024]. Prior research, in robotic timber assembly specifically, has explored various strategies to address these challenges, including the development of an external tracking system for end-effector pose correction in multi-robot assembly workflows Stadelmann et al. [2019], computational modeling of tolerance propagation Gandia et al. [2022], and adaptive, feedback-driven methods Helm et al. [2016], Adel et al. [2024], Ruan et al. [2023], Cote et al. [2024]. The successful and precise execution of timber joints often still depends on human intervention to correct deviations during nailing, gluing, and insertions Willmann et al. [2017], Adel et al. [2018], Adel [2020], Helmreich et al. [2022], Adel et al. [2024]. The challenges of uncertainty become even more pronounced in contact-sensitive assembly tasks, such as those for traditional timber joinery, where success depends not only on geometric precision but also on the ability to control excessive force interactions imposed by friction and tight tolerances. In such scenarios, even minor deviations introduced during fabrication or caused by material imperfections can easily lead to misalignment, resulting in jamming, stick-slip behavior, and material damage Apolinarska et al. [2021].
Handling force interactions in contact-rich robotic manipulation is often addressed through model-based impedance or admittance control laws Albu-Schäffer et al. [2007], Suárez-Ruiz et al. [2018], learning-based control policies Vecerik et al. [2019], Schoettler et al. [2020] or hybrid approaches combining both Johannsmeier et al. [2019], Hou et al. [2025]. In the context of assembly with timber joinery, a few studies have explored learning-based methods; Apolinarska et al. applied reinforcement learning for the motion control of an industrial robot to assemble single- and double-lap joints using robot pose and force/torque data from a wrist-mounted sensor Apolinarska et al. [2021]. For the single-lap tasks, they also investigated the policy robustness by introducing angular and linear offsets to the initial robot pose, grasping the top piece. Their results showed strong performance in zero-offset cases and revealed challenges of policy adaptation when angular and linear offsets, as well as angled or double insertions, were introduced. The findings underscored the importance of robust policy learning for handling geometric variations and positional deviations in contact-rich assembly tasks. Kramberger et al. also proposed a human–robot collaborative workflow using a learning-by-demonstration method that integrated a compliance controller into Cartesian-space dynamic movement primitives for perpendicular single-lap joint insertion; however, they relied on significantly loose tolerances Kramberger et al. [2022].
Other studies have explored semi-structured or open-loop methods for timber joinery assembly; Robeller et al. introduced a custom end-effector that generated vibration combined with a manually applied pulse force using a mallet to facilitate the connection of timber panels using dovetail joints Robeller et al. [2017]. Leung et al. demonstrated the use of an industrial robot in combination with manually placed distributed robotic clamps to assemble tight-fitting half-lap joints, highlighting the potential for precise and repeatable assembly. During their experiments, they observed challenges such as misalignments and collisions—largely arising from fabrication inaccuracies and deviations between digital and as-built models—that are common in construction-scale robotic applications Leung et al. [2021]. Rogeau employed visual feedback and fiducial markers for the assembly of wood panels with multiple mortise and tenon joints and reported successful full insertions in about 50% of their test cases across various joint configurations such as mortise tightness and tenon chamfer angles Rogeau [2023].
Model-based control can be particularly difficult when handling complex contact dynamics, such as those found in multi-body and multi-surface interactions or in the manipulation of deformable objects. To overcome the challenges of model-based control, sensory–motor policy learning methods such as imitation learning, meta learning, self-supervised learning, and reinforcement learning are increasingly employed as effective alternatives Levine et al. [2016], Finn et al. [2017], Levine et al. [2018], Kalashnikov et al. [2018], Brohan et al. [2023]. These policies are trained to condition on sensor observations such as images and force feedback to produce robot actions. Among the learning methods mentioned, behavior cloning (a subcategory of imitation learning) is widely adopted, as it replaces task-specific robot programming with human demonstrations, typically collected through teleoperation Seo et al. [2023], Wang et al. [2023], Shaw et al. [2023]. Policy learning from human demonstrations can be perceived as a sequenced supervised learning problem that maps a history of observations to the sequence of actions (robot commands). With access to sufficient task demonstrations, behavior cloning has shown promise in tasks involving challenging conditions such as deformable objects and bimanual coordination Zhao et al. [2023, 2025].
Diffusion policies Chi et al. [2024a], Ze et al. [2024], Zhao et al. [2025] represent a recent class of behavior cloning methods built upon Denoising Diffusion Probabilistic Models (DDPMs) Ho et al. [2020], Sohl-Dickstein et al. [2015] and Denoising Diffusion Implicit Models (DDIMs) Song et al. [2022]. DDPMs are generative models originally developed for high-quality image synthesis, transforming random noise into samples from a target distribution through an iterative denoising process. DDIMs extend this framework by formulating a non-Markovian reverse process, which allows the number of denoising steps used during inference to be reduced without changing the training procedure, thereby enabling significantly faster sampling. In the diffusion policy via robot action diffusion Chi et al. [2024a], the denoising process generates a sequence of robot trajectories at each inference step. The approach is known to naturally capture multimodal action distributions, meaning it can represent multiple distinct yet equally valid action sequences for completing a task, as often occurs in human demonstrations. The approach also maintains temporal coherence, meaning the policy predicts an entire sequence of future actions jointly rather than one step at a time. This joint prediction ensures that consecutive actions remain consistent (e.g., following one continuous path) rather than switching erratically between different predicted single actions, which can lead to unstable or jittery motion. Moreover, the method avoids the need for action discretization (breaking continuous movements into a fixed set of bins) or negative sampling (generating incorrect actions during training to guide learning) Chi et al. [2024a]. This removes common limitations of explicit behavior cloning methods Florence et al. [2020], Shafiullah et al. [2022], which can make it difficult to achieve high-precision actions. Diffusion policies also offer improved training stability compared to implicit behavior cloning methods Florence et al. [2022], Jarrett et al. [2020]. In addition, they have demonstrated robustness to perturbations, visual occlusions, and idle actions in some manipulation tasks Chi et al. [2024a].
Planning and control through contact remains a major challenge for both model- and learning-based approaches, due to the inherent complexity of contact dynamics Yang et al. [2025]. While diffusion policies have achieved notable success in various dexterous manipulation tasks, their effectiveness in contact-rich scenarios is still under active investigation Hou et al. [2025], Kang et al. [2025], Wu et al. [2025]. Additionally, despite significant progress in construction robotic research, the application of robot learning for planning and control, particularly for contact-sensitive tasks, remains underexplored. In this study, we evaluate the performance and robustness of diffusion policies through a case study on the assembly of timber mortise and tenon joints under fabrication uncertainty. The following section outlines the objectives, novelty, and contributions of this research.
1.2 Objectives and contributions
In the research presented in this paper, we develop a sensory–motor diffusion policy learning framework via behavior cloning for the contact-rich robotic manipulation at construction scale. The study pursues three primary objectives: (1) Demonstrate applicability and performance of diffusion policies through a case study on the robotic assembly of mortise and tenon timber joints; (2) Investigate policy robustness under construction-relevant uncertainties, simulated as randomized perturbations to the mortise position; and (3) Identify actionable guidelines for data collection strategies and policy parameter tuning that yield the most successful and robust performance, providing a practical reference for future research and deployment.
This paper presents one of the first systematic investigations in the AEC domain applying diffusion policy learning to robot planning and control. By enabling autonomous timber joinery assembly without reliance on hammering or human intervention, the proposed framework addresses physically demanding and craftsmanship-intensive tasks, offering a scalable learning framework for contact-sensitive manipulation tasks. The approach demonstrates potential for generalizing a wide spectrum of contact-rich assembly challenges, including complex timber joints, pipe fitting, and light metal framing. Ultimately, this work contributes to advancing robotic construction under uncertainty, bridging state-of-the-art generative policy learning with real-world AEC workflows, and provides both theoretical and practical advances that can accelerate adoption across the field.
2 Methods
This section outlines the experimental setup, data collection process, policy learning approach, and evaluation procedure used in our robotic manipulation system for timber joinery. We implement diffusion policy learning via behavioral cloning using robot pose and F/T data collected from teleoperated demonstrations.
Figure 2: Multi-robot setup consisting of two 6-axis industrial robotic arms.
2.1 Experimental setup
Our experimental setup consists of two primary systems for conducting this research. The first is a multi-robot setup equipped with customized end effectors for executing the timber assembly task. The second is a teleoperation system for manually operating the robotic setup and collecting data for demonstrations. This subsection details the specific hardware used for our experiments and their software interfaces.
The multi-robot setup consists of two 6-axis industrial robotic arms [^1] ABB IRB 4600 ABB Group , each with a 40 kg payload and a 2.55 m reach mounted on linear tracks (Fig. 2). Both robotic arms utilize custom pneumatically-driven end effectors designed for gripping long timber elements with varying cross-sectional profiles, as typically encountered in timber frame construction. One arm grasps the timber mortise in place and can systematically simulate construction deviations by adding positional offsets. The other arm grasps the timber tenon and is the primary agent for manipulation and completing the assembly task.
Because the industrial robotic arms do not provide native access to joint torque values, we developed the end effector on the tenon-grasping arm to include a 6-axis F/T sensor [^2] ATI Delta IP60, with SI-330-30 calibration ATI Industrial Automation to enable F/T data collection (Fig. 3). For safety, the end effector is also equipped with a pneumatic anti-collision sensor [^3] Schunk OPR 081-P00 Schunk , which both provides passive mechanical compliance and triggers a protective stop on the robotic arm under excessive force, safeguarding the robotic setup and F/T sensor. Motion control and pose feedback for the robotic arms utilize ABB’s Externally Guided Motion (EGM) feature ABB Robotics [2017], a UDP-based real-time control interface. F/T data is transferred over EtherCAT to a central Programmable Logic Controller (PLC) [^4] Beckhoff CX2062 Beckoff . All data communications are centralized on a desktop computer employing ROS [^5] ROS 2 Jazzy Jalisco Open Robotics [2023] interfaces and Python wrappers.
Figure 3: Gripper for holding the tenon piece equipped with anti-collision and F/T sensors.
The teleoperation system (Fig. 4) consists of a virtual reality (VR) system [^6] HTC VIVE Pro 2 HTC Corporation , and we developed a unilateral pose-to-pose transformation from the VR controller to the robotic arm’s end effector. This transformation is essential for aligning the coordinate frames of the VR system and the robotic setup, and for compensating for the human demonstrator’s orientation relative to the robot. For instance, when the demonstrator is facing the robotic arm, pushing the controller forward should consistently move the end effector away from the demonstrator, regardless of their position. In addition, this transformation includes an adjustable scaling parameter that maps the demonstrator’s control motions to proportionally smaller movements on the robot, enhancing the demonstrator’s fine motor control capabilities during high-precision tasks such as timber joint assembly. During teleoperation, robot motion and data collection (see Section 2.2) are enabled by the hand controller’s trigger. Similar to clicking and dragging a mouse, holding the trigger moves the robot in the desired direction, and releasing it allows the controller to be repositioned without affecting the robot or recording idle data. This allows for long motions to be performed in smaller, controlled segments. The VR controller interface utilizes OpenVR Valve Corporation [2024] in Python and ROS, integrating directly with the multi-robot setup by streaming mapped poses to EGM for fluid, real-time teleoperation.
Figure 4: VR-based teleoperation system for collecting human expert demonstrations.
2.2 Data collection
To train our diffusion policies, we created datasets consisting of expert demonstrations. An overview of the data collection workflow is illustrated in Fig. 5. During data collection, a human demonstrator utilized the teleoperation system to manipulate a vertical tenon into the mortise. The demonstrator initiates the task by simultaneously straightening the tenon and executing a downward motion to guide it into the mortise. To improve the robustness of the learned policy under fabrication uncertainties, we dedicated approximately 25% of the demonstrations to error recovery scenarios. In these cases, the demonstrator intentionally initiates contact with the edges of the mortise and performs a sliding maneuver along the surface to realign and insert the tenon, mimicking realistic recovery behaviors in the presence of misalignment. Each collected demonstration consists of timestamped robot poses and F/T data during motion (while the VR controller’s trigger is held).
Figure 5: Data collection and policy rollout workflow.
The raw data for each demonstration is recorded in real-time for each sensor interface: pose data is collected from EGM as a seven-dimensional (7D) vector (3D position in millimeters + 4D quaternion rotation of the tenon gripper end effector) every 12 ms (approximately 83 Hz), while F/T data is collected as a 6D vector (3D force in newtons + 3D torque in newton-meters) at 64 Hz after passing through an infinite impulse response (IIR) low-pass filter to attenuate frequency components above 64 Hz. This filtering step mitigates aliasing and improves signal quality by removing sensor noise and dynamic vibrations introduced by high-speed motion or structural resonances. To temporally align the different sample rates to a common time base, the pose and F/T data are interpolated to 60 Hz; position and F/T data are linearly interpolated, while rotation data uses spherical linear interpolation to preserve smooth interpolation.
Finally, each episode is post-processed to reduce unwanted noise in the demonstrations and remove periods of inactivity or unintentional stillness. Removing idle or near-idle periods is critical as it helps the policy training to focus on active, meaningful interactions rather than noise or moments where no significant movement occurs. We filter the pose and F/T data using a low-pass Butterworth filter Butterworth [1930], with filter parameters selected to reduce unwanted high-frequency noise while preserving the underlying movement patterns (Fig. 6). For the pose data, we used a 4th-order low-pass Butterworth filter with a 1 Hz cutoff, which provides a steeper roll-off and ensures that only low-frequency components characteristic of slower, continuous motion trajectories remain. In contrast, we used a 1st-order low-pass Butterworth filter with a higher 10 Hz cutoff to preserve sharper transitions (e.g., at the moment of contact) while effectively removing high-frequency noise that may come from mechanical oscillations or sensor inaccuracies. Actions are then derived from finite differences between successive pose data points, and any resulting vectors with magnitudes below a set threshold are removed to discard idle or near-idle actions.
Figure 6: Data filtering using a low-pass Butterworth filter. The visualized trajectory is for an error recovery scenario, demonstrating insertion after a collision with the mortise surface. Only the first 3 dimensions for each pose and F/T data are shown.
2.3 Policy learning
Our research employs CNN-based diffusion policy learning via behavioral cloning Chi et al. [2024a]. The method is based on DDPMs Sohl-Dickstein et al. [2015], Ho et al. [2020] and DDIMs Song et al. [2022], using DDPM for the forward noising process during training and DDIM for the reverse denoising process and faster inference during rollouts. Our diffusion policies predict a sequence of robot actions (i.e., the robot pose) conditional on the observation horizon of robot pose and F/T data. The foundation for our diffusion models is built upon the work of Bogert et al. van den Bogert et al. [2024].
When loading the dataset for training our diffusion policies, trajectories of 64 points are subsampled from each demonstration. This subsampling normalizes the trajectory length for each demonstration, as well as enables faster training iterations. Next, we converted pose quaternions to a continuous 6D rotation representation (by first converting to rotation matrices then concatenating the first two columns) to improve training stability Zhou et al. [2019], Chi et al. [2024a], resulting in 9D pose vectors. We then applied min-max normalization to scale all features to the range [-1, 1]. Given a pose observation horizon $T_{o}^{p}$, F/T observation horizon $T_{o}^{f}$, and action prediction horizon $T_{p}$, each training sample at timestep $t$ consists of an observation feature $\mathbf{O}_{t}$ concatenating the last $T_{o}^{p}$ and $T_{o}^{f}$ and a corresponding action sequence $\mathbf{A}_{t}$ consisting of the next $T_{p}$ absolute poses.
During the rollouts, $\mathbf{A}_{t}$ is predicted conditioned on $\mathbf{O}_{t}$. The reverse denoising process for action prediction starts from ${A}^{K}$ sampled from Gaussian noise. After iterating $K$ denoising steps, we output a desired noise-free action ${A}^{0}$ at step 0. The action at each reverse step k1k-1 is calculated as Chi et al. [2024a]:
$\mathbf{A}_{t}^{k-1}=\alpha\left[\mathbf{A}_{t}^{k}-\gamma\epsilon_{\theta}+\mathcal{N}\left(0,\sigma^{2}I\right)\right]$ (1)
where $\epsilon_{\theta}$ is the learned predicted noise and $\mathcal{N}(0,\sigma^{2}I)$ is Gaussian noise. The $\alpha$, $\gamma$, $\sigma$ are functions of the iteration step $k$ and are defined through a selected noise schedule. In our case, a squared cosine noise schedule Nichol and Dhariwal [2021] is used. The noise prediction network $\epsilon_{\theta}$ is learned using a 1D CNN U-Net architecture by minimizing the following mean squared error (MSE) loss Chi et al. [2024a]:
$\mathcal{L}=\sum\left\lVert\epsilon^{k}-\epsilon_{\theta}\left(\mathbf{O}_{t},\mathbf{A}^{k}_{t},k\right)\right\rVert$ (2)
where $\epsilon^{k}$ is the sampled random noise with known variance at iteration $k$. The predicted noise $\epsilon_{\theta}(\mathbf{O}_{t},\mathbf{A}^{k}_{t},k)$ is a function of $\mathbf{O}_{t}$, the pose and F/T observation feedback preceding time $t$, $\mathbf{A}^{k}_{t}$ = $\mathbf{A}^{0}+\epsilon^{k}$ with $\mathbf{A}^{0}$ sampled from starting noise-free actions, and iteration step $k$. As implemented in Chi et al. [2024a], we add the sampled random noise $\epsilon^{k}$ to the noise-free actions $\mathbf{A}_{0}$ in $K$ forward steps (original DDPM method).
Policy training was conducted on high-performance computing (HPC) clusters [^7] Princeton Research Computing’s Della cluster, which provide powerful computational resources for efficiently processing large datasets and enabling faster, scalable learning by leveraging parallel computing capabilities. Each policy was trained for a fixed amount of time (1 hour), with each training lasting at least 1,000 epochs. Only the model weights corresponding to the epoch with the lowest validation loss were used to evaluate performance during rollouts. To account for training variance, we trained 4 models for each set of parameters and averaged their success rates when computing their performance (see Section 2.4).
2.4 Evaluation
To evaluate the performance of our trained policies, we perform rollouts on the experimental setup (Section 2.1). The starting position for the tenon piece during all rollouts is fixed and identical to the demonstrations. During policy rollouts, we specify additional parameters for the number of inference steps $K_{\text{inf}}$, DDIM/DDPM interpolation $\eta$ Song et al. [2022], and action execution steps $T_{a}$. The number of inference steps ($K_{\text{inf}}\leq K$) controls the number of steps in the reverse denoising process during policy inference. Decreasing $K_{\text{inf}}$ reduces computation and increases rollout speed, which is important for real-time policy execution. However, smaller $K_{\text{inf}}$ may also reduce the quality of noise sampling or action precision.
The amount of stochasticity during the reverse denoising process is controlled by the interpolation parameter $\eta\in[0,1]$, where $\eta=0$ is the DDIM process and $\eta=1$ is the original DDPM process. This means that when $\eta=0$, the standard deviation of the noise added during sampling is 0, meaning that the reverse process is fully deterministic, producing high-fidelity imitations of demonstrations, and when $\eta=1$, the reverse process is fully stochastic, introducing broader behavioral coverage with exploration and variability.
The number of action execution steps $T_{a}\leq T_{p}$ denotes that of the $T_{p}$ predicted actions at each time step; we execute the first $T_{a}$ actions. Decreasing the value of $T_{a}$ increases the frequency of replanning over a receding horizon, increasing the responsiveness of the system; however, it can also potentially decrease the speed of rollouts by spending more time on inference. For values of $T_{a}>2$, we apply an average pooling with a kernel size of $T_{a}-1$ across each dimension of the actions, smoothing the action sequence.
In contrast to data collection during demonstrations, observations during rollouts are not temporally aligned and instead utilize the latest available pose and F/T data when the observation is retrieved. These observations are normalized using the same normalization parameters as the dataset before being passed to the policy for inference. The output action sequence is similarly denormalized back into the original scale using the same action normalization parameters during training. These predicted actions are then sent through EGM to the robot controller, and we use a similar communication protocol to the data collection to receive pose and F/T feedback (see Fig. 5).
The policy performance is determined by rolling out 5 times for each of the 4 trained models with fixed parameters, to account for stochasticity during test time, resulting in an average success rate (Avg. SR) across 20 combined rollouts. A rollout is recorded as a success if the tenon element fully inserts into the mortise hole without triggering any collision errors (i.e., excessive force at either the collision sensor or the robot joints). A rollout is recorded as a failure if any collision errors are triggered by the industrial robot controller or if the policy does not appear to be making any meaningful progress towards task completion, as determined by the operator.
3 Experiments
We conducted a two-phase study: in Phase 1 we performed static experiments with a fixed mortise location to establish a general performance evaluation for the diffusion policy in our industrial setup; in Phase 2 we performed uncertainty experiments with randomized hole offsets to evaluate the robustness of the diffusion policies.
The dimensions of the mortise and tenon we used for the experiments are illustrated in Fig. 7. The tolerance between the tenon cheeks and the mortise walls was 0.1 mm, and the tolerance between the tenon edges and the mortise walls was about 2 mm. The tight tolerance of the tenon cheeks is what ensures a structural fit for a mortise and tenon joint. This figure also visualizes the boundary of the 5 mm and 10 mm offsets used in the uncertainty experiments (see Section 3.2). The tenon is aligned along the world’s z-axis, while the mortise piece lies horizontally along the x-axis. Each demonstration and rollout begins with the tenon element grasped at a fixed starting pose, positioned approximately 15 mm vertically above the mortise hole with an approximately 6°angular deviation about the y-axis.
Figure 7: Mortise and tenon terminology, dimensions, and reference frame.
3.1 Phase 1: Static experiments
The objective of the Phase 1 experiments is to show the application and performance of diffusion policies in a building-scale robotic assembly setting. For this experiment, we created a dataset of 100 demonstrations following the data collection method described in Section 2.2, with the exception that no demonstrations included error recovery cases and demonstrations were not post-processed. To simplify the assembly task, the mortise position was held fixed across all demonstrations and during rollouts. Through repeated experimentation, we identified the best-performing set of parameters (reported in Section 4.1), which were then used to initialize parameter tuning in the Phase 2 experiments described in the following subsection.
3.2 Phase 2: Uncertainty experiments
The objective of the Phase 2 experiments is to demonstrate the robustness of diffusion policies in handling fabrication uncertainties for contact-rich timber joinery. More specifically, we evaluate how the policies perform when the mortise is subjected to random position offsets simulating fabrication tolerances.
We created a dataset of 400 demonstrations following the data collection method outlined in Section 2.2. To simulate fabrication tolerances, the position of the mortise in these demonstrations was perturbed by a displacement vector sampled in polar coordinates. The angular component was drawn uniformly from $0,2\pi)$ radians, and the radial component was drawn uniformly from $[0,10]$ mm. Unlike uniform sampling over the area of a disk, this approach yields a higher density of samples near the origin, biasing the distribution toward smaller offsets while still allowing for the possibility of larger deviations.
During rollouts, we evaluated each policy at three goal offset distances: 0 mm (no offset), 5 mm, and 10 mm. These values were selected to evaluate the policies under no uncertainty (0 mm), uncertainty within the demonstration distribution (5 mm), and uncertainty at the edge of the demonstration distribution (10 mm). For each offset condition, the goal position was randomized uniformly along the circumference of a circle with a radius of the specified distance. The success rate of a policy was computed separately for each offset distance, and its overall performance was reported as the average success rate across all three conditions.
In the first part of the Phase 2 experiments, we systematically tuned the parameters of the diffusion policy to maximize the success rate. Starting from a candidate configuration derived from the Phase 1 experiments (Section 3.1), we performed sequential, greedy parameter tuning, optimizing one parameter at a time while holding others fixed. The resulting best-performing policy is referred to as the full model. The impact of individual parameters on policy performance is reported in Section 4.2.1.
In the second part of the Phase 2 experiments, we conducted two ablation studies to assess the impact of F/T observations and demonstration count on policy performance. In the F/T ablation study, we evaluated the performance of the full model with F/T inputs masked (i.e., input F/T values are set to 0), as well as a variant trained using only pose observations. For the demonstration ablation, we trained policies using 25, 50, 100, and 200 demonstrations, while keeping all other parameters identical to the full model. The results of these studies are presented in Sections 4.2.2 and 4.2.3.
4 Results and discussion
Across all experiments, policy training exhibited consistent and stable convergence. Training and validation MSE losses remained closely aligned throughout policy training, suggesting minimal overfitting. Both losses decreased steadily over the first 300 epochs, followed by a gradual plateau beginning at approximately 600 epochs. An example of the MSE loss curves is illustrated in Fig. 8 for the full model (see Section 4.2.1). The loss curves were first averaged across all 4 training iterations, then smoothed using a Gaussian filter with a standard deviation $\sigma=10$ for visualization.
Figure 8: Smoothed mean squared error (MSE) training and validation loss curves for Policy 4, averaged then smoothed across all training iterations. The shaded area represents $\pm 1$ standard deviation.
Fig. 9 illustrates two example rollout trajectories: the top sequence shows a successful rollout where the tenon was first reoriented and then smoothly inserted into the mortise; the bottom sequence shows an unsuccessful rollout where the tenon did not finish reorienting before hitting the edge of the mortise and was unable to recover correctly.
Figure 9: Example of an insertion sequence for a successful rollout (top row) and an unsuccessful rollout (bottom row).
4.1 Phase 1: Static experiments
After repeated testing and parameter tuning, the policy achieved a 100% average success rate over 20 rollouts using the parameter values listed in Table 1. This performance demonstrates the capability of diffusion policies to reliably assemble a timber mortise and tenon joint under no uncertainty. These parameter values were subsequently used as the initialization point for parameter tuning in the Phase 2 experiments.
Table 1: Parameters for the Phase 1 experiments.
Symbol Description Value
$L_{r}$ Learning rate $1\mathrm{e}{-5}$
$W_{d}$ Weight decay rate $1\mathrm{e}{-6}$
$T^{p}_{o}$ Pose observation horizon 1
$T^{f}_{o}$ Force observation horizon 1
$T_{p}$ Action prediction horizon 8
$T_{a}$ Action execution steps 4
$K$ Forward noising steps 128
$K_{\text{inf}}$ Inference/denoising steps 32
$\eta$ DDIM/DDPM interpolation 0.5
During the first phase of the experiments, we also evaluated encoder networks for pose and F/T inputs to the diffusion policy. Each encoder consisted of a single fully connected layer that projected the raw input vector to a latent representation of fixed length (e.g., 64 or 128 units). The aim was to learn a compact feature embedding prior to policy integration. However, policies incorporating these encoders underperformed relative to those that directly concatenated the unprocessed pose and F/T observations. Future research may explore the impact of more sophisticated encoders, such as multi-layer, frequency-based, or attention-based architectures, on policy performance Lee et al. [2020], Hou et al. [2025].
The Phase 1 experiments also revealed several practical challenges that we addressed during the Phase 2 experiments. First, as the human demonstrator became increasingly familiar with the setup, the demonstrations tended to become faster and more fluid over time. Although random sampling during dataset construction mitigated this effect, the demonstrations lacked examples of mistakes and error recovery for the policy to learn from. This limitation was later addressed in our revised data collection procedure (Section 2.2), in which a specified percentage of demonstrations was deliberately dedicated to error recovery scenarios.
A second challenge arose from the gradual physical degradation of the experimental setup as a consequence of the large construction-scale forces repeatedly exerted on the timber mortise and tenon. This degradation manifested in several ways: wear on contact surfaces (increasing tolerance and making insertion easier), small chips breaking off the tenon edge (slightly decreasing tolerance due to glue reattachment), and minor shifts of the mortise and tenon within the gripper grasp. Such occurrences are expected in real-world construction environments, but they also introduce additional sources of uncertainty. Future research should explicitly account for these fabrication-induced variations when developing and evaluating assembly policies.
4.2 Phase 2: Uncertainty experiments
4.2.1 Full model parameter tuning
The sequential parameter tuning in Phase 2 was limited to three stages: (1) pose and F/T observation horizons; (2) action prediction horizon; (3) number of inference steps and DDIM/DDPM interpolation. The learning rate, weight decay, and forward noising steps were fixed at the values determined in Phase 1.
Table 2 summarizes the results from the first two stages. The highest average total success rate of 75%, computed over the three mortise offset conditions, was achieved using an observation horizon of 1 for both pose and F/T modalities combined with a longer action prediction horizon of 16. This configuration suggests that, for the uncertainty conditions tested, the policy benefited from focusing on the most recent state information while planning actions further into the future, potentially allowing it to better anticipate and correct for deviations introduced by positional offsets.
Table 2: Success rates for training parameter tuning.
Policy # Demos $T_{p}$ $T^{p}_{o}$ $T^{f}_{o}$ $K_{\text{inf}}$ $\eta$ $T_{a}$ D (mm) Avg. SR (%) Avg. total SR (%)
[…]
5 55
10 65
2 400 8 2 1 32 0.5 4 0 85 63
5 65
10 40
3 400 8 2 2 32 0.5 4 0 80 60
5 60
10 40
[…]
5 65
10 60
  • D = randomized hole offset.
Table 3 summarizes the final stage of rollout parameter tuning, evaluating different combinations of inference steps and DDIM/DDPM interpolation values. These tests were conducted only for mortise offsets of 5 mm. The initial parameter values selected from the Phase 1 experiments achieved the highest performance, with a 65% average success rate, substantially outperforming other tested configurations.
Table 3: Success rates for rollout parameter tuning.
Policy # Demos $T_{p}$ $T^{p}_{o}$ $T^{f}_{o}$ $T_{a}$ D (mm) $K_{\text{inf}}$ $\eta$ Avg. SR (%)
[…]
16 0.5 35
64 0.5 40
32 1 35
32 0.25 50
  • D = randomized hole offset.
4.2.2 Force/torque ablation study
Table 4 summarizes the results of the F/T ablation study, and Fig. 10 illustrates the contribution of F/T feedback to task performance. As expected, both ablated policies underperformed relative to the full model. However, a notable pattern emerged when comparing the pose-only policy to the masked full model. When F/T inputs were masked in the full model, the tenon consistently collided with the mortise, reflecting the model’s inability to detect contact without force feedback. Surprisingly, the pose-only policy occasionally succeeded under the 10 mm mortise offset condition, despite lacking force information. The large standard error in model variance (Fig. 10) for this condition suggests that these successes were likely the result of random factors rather than reliable inference. This highlights the critical role of force feedback in consistent policy performance, particularly under larger fabrication uncertainties.
Table 4: Success rates for the F/T ablation study.
Policy Modality # Demos $T_{p}$ $T^{p}_{o}$ $T^{f}_{o}$ $K_{\text{inf}}$ $\eta$ $T_{a}$ D (mm) Avg. SR (%) Avg. total SR (%)
4 PF 400 16 1 1 32 0.5 8 0 100 75
5 65
10 60
4 PF* 400 16 1 1 32 0.5 8 0 70 45
5 45
10 20
9 P 400 16 1 1 32 0.5 8 0 85 53
5 35
10 40
  • D = randomized hole offset; PF = models trained with pose and force data; P = models trained with pose-only data.
  • *The forces were masked during rollouts.
Figure 10: Success rates of diffusion policies with force masked (PF*) and trained only on pose data (P) compared to the full model (PF), evaluated at three mortise offsets (0, 5, and 10 mm). Error bars indicate the standard error of the mean (SEM) computed across the 4 independently trained models for each set of parameters.
4.2.3 Demonstration quantity ablation study
Table 5 summarizes the results of the demonstration quantity ablation study. Policies trained on only 25 demonstrations failed to fully converge and achieved no success during rollouts, and were therefore excluded from the reported results. Fig. 11 shows a pronounced drop in success rate at the 10 mm mortise offset when the number of demonstrations was reduced, likely due to insufficient coverage of the state-action space in the demonstration data. The results suggest a performance inflection point between 200 and 400 demonstrations, beyond which the policy gains a substantial improvement in its ability to handle positional uncertainty. This trend indicates that under higher uncertainty, diffusion policies may require a critical mass of demonstrations to adequately represent recovery behaviors.
Notably, the total number of demonstrations required to achieve robust performance was much higher than anticipated. This is likely attributable to the trajectory subsampling strategy used during policy learning (Section 2.3), which significantly reduced the total number of training samples available to the policy. Future research should utilize alternative data preprocessing strategies to maintain a sufficient quantity of training data to enable reliable policy learning.
Table 5: Success rates for the demonstration quantity ablation study.
Policy # Demos $T_{p}$ $T^{p}_{o}$ $T^{f}_{o}$ $K_{\text{inf}}$ $\eta$ $T_{a}$ D (mm) Avg. SR (%) Avg. total SR (%)
4 400 16 1 1 32 0.5 8 0 100 75
5 65
10 60
5 200 16 1 1 32 0.5 8 0 80 60
5 75
10 25
6 100 16 1 1 32 0.5 8 0 80 58
5 75
10 20
7 50 16 1 1 32 0.5 8 0 70 48
5 55
10 20
  • D = randomized hole offset.
Figure 11: Success rates of diffusion policies trained with different numbers of demonstrations (50, 100, 200, and 400) evaluated at three mortise offsets (0, 5, and 10 mm). Error bars indicate the standard error of the mean (SEM) computed across the 4 independently trained models for each set of parameters.
5 Conclusion
The experimental results presented in this work demonstrate that diffusion policies can achieve stable convergence and high success rates in a building-scale contact-rich assembly task representative of real-world construction. In Phase 1, we established a baseline by identifying parameter settings capable of achieving 100% average success rate in a deterministic mortise and tenon joint assembly task. In Phase 2, we evaluated policy robustness under controlled fabrication uncertainties, identifying parameter configurations that improved generalization to offsets of the mortise position, quantifying the critical role of force/torque feedback, and determining the demonstration quantity threshold necessary for reliable performance.
Together, these findings address the study’s core objectives: they validate the applicability of diffusion policies to building-scale robotic assembly and systematically characterize their robustness under fabrication uncertainty. Beyond the performative metrics, the experiments also yielded actionable insights for parameter tuning, sensor fusion, and demonstration count, forming a generalizable methodology for future contact-rich construction and assembly tasks. By demonstrating that diffusion policies can be systemically adapted and tuned to handle fabrication uncertainties without manual intervention, this research establishes a scalable, feedback-driven robot learning framework that addresses not only the technical demands of precision assembly, but also broader industry challenges such as labor shortages and the need to reduce reliance on physically taxing, repetitive, and high-skill manual work. Beyond timber joinery, the presented methodology can be generalized to a wide range of construction and manufacturing scenarios, from complex timber connections and pipe fitting to light metal framing, where expert craftsmanship has traditionally been essential. In doing so, this work advances robotic construction under uncertainty while contributing to safer and more efficient building practices, positioning robots as capable collaborators in the evolving construction workforce.
5.1 Limitations and future work
While this work demonstrates the feasibility and robustness of applying sensory–motor diffusion policy learning to contact-rich robotic assembly at construction scale, several limitations remain; our approach relies exclusively on robot pose and force data to train the policy and evaluate robustness, whereas diffusion policies for robot planning and control are commonly formulated as visuomotor policies using end-to-end mappings from RGB or tactile images to robot commands Chi et al. [2024a, b]. Our simplification of the sensory input space increases training speed but may limit the model’s ability to generalize to visually complex or occlusion-prone construction scenarios, such as joint assembly in cluttered framing environments or alignment tasks in multi-material panel installations. Moreover, incorporating expressive state-of-the-art learning models, such as diffusion transformers or vision-language-action (VLA) models Octo Model Team et al. [2024], Kim et al. [2024], could enhance representation learning, enabling better generalization, scalability, and transfer across tasks. These capabilities are particularly relevant for building-scale construction where robots must integrate multiple sensing modalities and operate across varied materials, such as combining timber, metal fasteners, and composite claddings in a single assembly sequence.
Furthermore, a second limitation stems from our reliance on single-task behavior cloning from human demonstrations, which introduces both practical and algorithmic challenges. Collecting high-quality demonstrations and conducting real-world rollouts is time-intensive and can cause material degradation, especially in contact-sensitive scenarios or fragile materials like softwoods, architectural veneers, or insulation panels. Moreover, the learned policy is inherently bounded by the quality and diversity of human performance, raising the question of how to scale beyond demonstrator capabilities. This is especially critical for more dexterous and varied AEC tasks, such as assembling shingle patterns on curved surfaces, handling flexible ductwork, or manipulating deformable architectural fabrics Craney and Adel [2020], Graser et al. [2021]. Future work could address these limitations by exploring hybrid learning paradigms that combine demonstration-based training with simulation pretraining, self-improvement cycles, or reinforcement learning, enabling policies to acquire physical reasoning skills and recover from unseen conditions TRI LBM Team et al. [2025]. Additionally, emerging research on general-purpose robot foundation models, such as visuomotor foundation models and Large Behavior Models (LBMs) TRI LBM Team et al. [2025], Zhao et al. [2023], presents a promising direction for learning transferable policies across a broad range of dexterous manipulation tasks.
Acknowledgments
This research was supported by the National Science Foundation (NSF, Award No. 2122271), Princeton Catalysis Initiative, and the School of Architecture (SoA) at Princeton University. The authors would like to thank Ruxin Xie, Zhengyi Chen, and Roman Ibrahimov at the Adel Research Group (ARG) and SoA for their invaluable support in hardware development and data collection.
Data availability statement
The data used in this study are available upon request.

### v2

@@
 Abstract
-Construction uncertainties such as fabrication inaccuracies and material imperfections pose a significant challenge to contact-rich robotic manipulation by hindering precise and robust assembly.
-In this paper, we explore the performance and robustness of diffusion policy learning as a promising solution for contact-sensitive robotic assembly at construction scale, using timber mortise and tenon joints as a case study.
-A two-phase study is conducted: first, to evaluate policy performance and applicability; second, to assess robustness in handling fabrication uncertainties simulated as randomized perturbations to the mortise position.
-The best-performing policy achieved a total average success rate of 75% with perturbations up to 10 mm, including 100% success in unperturbed cases.
-The results demonstrate the potential of sensory–motor diffusion policies to generalize to a wide range of complex, contact-rich assembly tasks across construction and manufacturing, advancing robotic construction under uncertainty and contributing to safer, more efficient building practices.
+Fabrication uncertainties, such as tolerance accumulation and material imperfections, pose a significant challenge to contact-rich robotic manipulation in construction by hindering precise and robust assembly.
+In this paper, we investigate the performance and robustness of diffusion policy learning for contact-rich assembly at the construction scale, using a tight-fitting timber mortise and tenon joint as a case study.
+A two-phase experimental study is conducted: first, to evaluate baseline policy performance and applicability; second, to assess policy robustness under fabrication-induced uncertainties modeled as randomized perturbations to the mortise position.
+The diffusion policy is trained on teleoperated demonstrations using an industrial robotic arm conditioned on end-effector pose and force/torque feedback.
+The best-performing policy achieved a total average success rate of 75% under perturbations up to 10 mm, including 100% success in unperturbed cases.
+The results demonstrate the potential of sensory–motor diffusion policies to enable high-precision contact-rich manipulation on large-scale industrial robotic arms, reducing reliance on skilled manual intervention.
+This work advances robotic construction under uncertainty and provides practical insights for deploying learning-based control in real-world Architectural, Engineering, and Construction applications.
 keywords:
-Contact-rich manipulation, Robotic assembly, Fabrication uncertainty, Diffusion policy, Timber joinery, Robot learning, Construction robotics
+Contact-rich manipulation, Diffusion policy, Fabrication uncertainty, Timber joinery, Robotic assembly, Construction robotics
 1 Introduction
@@
 [2023].
-The industry often requires workers to perform repetitive and physically demanding tasks, such as lifting and positioning heavy components, which can contribute to chronic injuries and reduced long-term workforce capacity Laukkanen [1999], Arndt et al.
+The industry often requires workers to perform repetitive, physically demanding tasks, such as lifting and positioning heavy components, which can lead to chronic injuries and reduced long-term workforce capacity Laukkanen [1999], Arndt et al.
 [2005].
 Automation, and in particular robotic assembly, offers the potential to address these issues by enhancing precision, increasing efficiency, enabling mass customization, and reducing the physical burden on workers Bock [2015], Musarat et al.
-[2024], Chen and Adel [2025].
-Recent advances in construction robotics have increasingly harnessed the precision and programmability of industrial robotic arms to execute complex and customized assemblies that would be difficult to achieve through manual methods Graser et al.
-[2021], Chen and Adel [2025].
+[2024].
+Recent advances in construction robotics have increasingly utilized industrial robotic arms, which play a central role in large-scale robotic assembly due to their high payload capacity and long reach to execute customized and complex assemblies that would be difficult to achieve manually Chen and Adel [2025].
 Despite these advancements, translating the capabilities of industrial robots from manufacturing to the uncertain and unstructured conditions of construction, characterized by fabrication inaccuracies, material imperfections, as well as dynamic and evolving environments, remains a significant challenge, making it difficult to achieve the same levels of repeatability, quality, and precision that are standard in factory settings Adel et al.
@@
 [2017], Apolinarska [2018], Adel et al.
-[2018], Adel [2020, 2022], Chai et al.
+[2018], Adel [2020], Graser et al.
+[2021], Chai et al.
 [2022], Lauer et al.
-[2023], where many workflows rely on planar face-to-face or butt joints, whose geometric simplicity allows the use of pre-programmed trajectories for successful assembly.
-In contrast, the manipulation of traditional timber joinery such as mortise and tenon, lap, or scarf joints requires contact-sensitive insertion and interlocking, demanding high precision and control of contact forces Apolinarska [2018], Leung et al.
+[2023], where many workflows rely on planar face-to-face or butt joints, whose geometric simplicity allows the use of preprogrammed trajectories for successful assembly.
+In contrast, the manipulation of timber joinery such as mortise and tenon, lap, or scarf joints requires contact-rich insertion and interlocking, demanding high precision and control of contact forces Apolinarska [2018], Leung et al.
 [2021].
-Similar to many contact-rich robotic manipulation tasks Suomalainen et al.
-[2022], the assembly of these joints can lead to misalignment and jamming if not coupled with real-time control for robust error recovery Apolinarska et al.
+In these scenarios, misalignment and jamming can occur if insertion is not coupled with real-time control and robust error recovery Apolinarska et al.
 [2021], Kramberger et al.
 [2022].
-As a result, these contact-sensitive tasks are often delegated to skilled human workers, who intervene to clamp, hammer, and adjust fits, compensating for fabrication inaccuracies and material imperfections Yang et al.
+As a result, these tasks are often delegated to skilled human workers, who intervene to clamp, hammer, and adjust fits, compensating for fabrication inaccuracies and material imperfections Yang et al.
 [2024].
@@
 Furthermore, it increases reliance on a shrinking pool of skilled labor and entails physically demanding, labor-intensive work that can contribute to worker fatigue and injury.
-Figure 1: Overview of our method. A CNN-based diffusion policy is trained conditional on end-effector pose and F/T data observations to predict sequences of robot actions.
+Figure 1: Overview of our method. A CNN-based diffusion policy is trained conditional on end-effector pose (i.e., position and orientation) and F/T observations to predict sequences of robot actions.
 Timber joinery is a centuries-old construction technique of geometrically interlocking timber components to create durable structural connections with minimal reliance on fasteners or adhesives Benson [1981].
 These joints are valued for their enhanced structural performance, rotational stiffness, and resilience under seismic and fire conditions Fang and Mueller [2023], Fang [2021].
-While they offer significant structural advantages, their complex geometry and contact-rich assembly make them particularly sensitive to the uncertainties inherent to construction tasks, including fabrication inaccuracies from cut length and angle deviations and component shifting during picking, placing, and fastening, as well as material imperfections from dimensional variations and hygrothermal expansion Adel et al.
+However, their complex geometry and contact-dominated assembly process make them particularly sensitive to fabrication inaccuracies and material imperfections, including cut length and angle deviations, component shifting during picking, placing, and fastening, and dimensional changes due to hygrothermal effects Adel et al.
 [2024].
-Collectively, these factors further increase the complexity of automating the assembly of such joints in construction settings.
-Despite the significance of this challenge, automated assembly of traditional joinery remains nascent in robotic workflows within the Architecture, Engineering, and Construction (AEC) domain, particularly when accounting for fabrication and material uncertainty.
-Overcoming these limitations requires feedback-driven and closed-loop planning and control methods capable of adapting to unpredictable conditions inherent in construction workflows.
+These factors significantly complicate robotic automation of joinery assembly.
+Despite the significance of this challenge, automated assembly of traditional joinery remains nascent in robotic workflows within the Architecture, Engineering, and Construction (AEC) domain, particularly when explicitly accounting for fabrication-induced and material-driven variability.
+Overcoming these limitations requires feedback-driven, closed-loop planning and control methods capable of adapting to the unpredictable and evolving conditions inherent in construction workflows.
 Learning-based methods have increasingly replaced traditional control-based methods for planning robot actions.
-Sensory–motor policy learning methods integrate sensor observations, planning, and control into a single end-to-end system, originally implemented using deep convolutional neural networks (CNNs) Levine et al.
+Sensory–motor policy learning methods integrate sensor observations, planning, and control into a single end-to-end mapping, originally implemented using deep convolutional neural networks (CNNs) Levine et al.
 [2016].
-These methods enable the robot to learn how to act in response to what it senses without requiring separately engineered modules for perception, planning, state estimation, or low-level control.
-These learned control policies map raw observation inputs, such as force feedback and images captured from a camera, directly to motor commands or robot actions using various learning-based methods Finn et al.
+These learned control policies map raw observations, such as images captured by a camera and force feedback, directly to motor commands or robot actions using various learning-based methods Finn et al.
 [2017], Levine et al.
@@
 Among these methods, diffusion policy learning via behavior cloning Chi et al.
-[2024a] has shown outstanding performance for various manipulation scenarios Chi et al.
+[2024a] has demonstrated strong performance across various manipulation scenarios Chi et al.
 [2024b], Hou et al.
@@
 [2025].
-The key advantages of diffusion policies lie in their ability to learn multi-modal action distributions, capturing multiple plausible ways to perform a task, while maintaining training stability and predicting coherent sequences of future actions Chi et al.
+The key advantages of diffusion policies lie in their ability to learn multimodal action distributions that capture multiple plausible ways to perform a task, while maintaining training stability and predicting coherent sequences of future actions Chi et al.
 [2024a].
-The research presented in this paper tackles the challenge of contact-rich robotic manipulation under construction uncertainties, where fabrication inaccuracies and material imperfections hinder precise and robust assembly.
-We address this by implementing diffusion policy learning, trained on robot pose and force/torque (F/T) data from teleoperated demonstrations, to plan and control the construction-scale assembly of mortise and tenon joints as a case study (Fig.
+This paper addresses contact-rich robotic manipulation under fabrication uncertainty, defined here as the combined influence of tolerance accumulation across assembly steps and geometric variability arising from material imperfections, both of which introduce positional misalignment in tight-clearance timber joints and hinder precise and robust robotic assembly.
+We develop a sensory–motor diffusion policy learning framework via behavior cloning, trained on robot end effector pose and force/torque (F/T) data from teleoperated demonstrations, to plan and control the construction-scale assembly of mortise and tenon joints as a case study (Fig.
 1).
-1.1 Related work
+The study pursues three primary objectives: 1) Demonstrate the applicability and performance of diffusion policies through a case study on the robotic assembly of mortise and tenon timber joints with sub-millimeter joint clearance; 2) Investigate policy robustness under fabrication uncertainties, simulated as randomized perturbations to the mortise position; and 3) Identify actionable guidelines for data collection strategies and policy parameter tuning that yield the most successful and robust performance, providing a practical reference for future research and deployment on construction-scale industrial robotic arms.
+This paper presents one of the first systematic investigations in the AEC domain applying diffusion policy learning to robot planning and control.
+By enabling autonomous timber joinery assembly without reliance on human intervention, the proposed framework addresses physically demanding, craftsmanship-intensive tasks and offers a learning framework for high-precision contact-rich manipulation on large-scale industrial robots.
+The approach demonstrates its potential to generalize across a broad spectrum of contact-rich manipulation for assembly, including complex timber joints, pipe fittings, and light metal framing.
+Ultimately, this work advances robotic construction under uncertainty by bridging state-of-the-art sensory–motor policy learning with real-world AEC workflows and provides practical advances that can accelerate adoption across the field.
+2 Related work
+This section reviews prior work on robotic timber assembly, contact-rich manipulation, and sensory–motor policy learning, and learning-based approaches for construction applications.
+We first discuss the methods used for robotic timber assembly under uncertainty, including the contact-rich manipulation of timber joints.
+Next, we summarize recent advances in sensory–motor policy learning.
+Finally, we situate existing robot learning efforts in construction tasks, motivating the present study.
+2.1 Robotic timber assembly
 In construction with discrete elements, such as brickwork and timber or steel framing, uncertainty remains a fundamental challenge to achieving precise and robust robotic assembly.
-In timber framing, which serves as the case study in this research, these uncertainties arise from material imperfections (e.g., dimensional deviations, deformation, shrinkage/expansion due to moisture) and fabrication inaccuracies (e.g., cut length and angle errors, pose deviations from calibration or handling).
-If unaccounted for, these variabilities accumulate during assembly, leading to misalignments, part collisions, and robot failures when trajectories are executed directly from as-planned digital models  Adel et al.
+In robotic timber assembly, which serves as the case study in this paper, these uncertainties arise from material imperfections (e.g., dimensional deviations, deformation, shrinkage/expansion due to moisture) and fabrication inaccuracies (e.g., cut length and angle errors, robot pose or grasp pose deviation).
+If unaccounted for, these perturbations cause error accumulation during assembly, leading to misalignment, part collision, and task failure when trajectories are executed directly from as-planned digital models  Adel et al.
 [2024].
-Prior research, in robotic timber assembly specifically, has explored various strategies to address these challenges, including the development of an external tracking system for end-effector pose correction in multi-robot assembly workflows Stadelmann et al.
-[2019], computational modeling of tolerance propagation Gandia et al.
-[2022], and adaptive, feedback-driven methods Helm et al.
+Prior research in robotic timber assembly has explored various strategies to address these uncertainties, including computational modeling of tolerance propagation Gandia et al.
+[2022], adaptive, feedback-driven methods Helm et al.
 [2016], Adel et al.
@@
 [2023], Cote et al.
-[2024].
-The successful and precise execution of timber joints often still depends on human intervention to correct deviations during nailing, gluing, and insertions Willmann et al.
+[2024], and external tracking systems for end-effector pose correction in multi-robot assembly workflows Stadelmann et al.
+[2019], Adel [2020].
+However, the successful and precise execution of timber joints often still depends on human intervention to correct deviations during nailing, gluing, and insertions Willmann et al.
 [2017], Adel et al.
 [2018], Adel [2020], Helmreich et al.
-[2022], Adel et al.
 [2022].
-The challenges of uncertainty become even more pronounced in contact-sensitive assembly tasks, such as those for traditional timber joinery, where success depends not only on geometric precision but also on the ability to control excessive force interactions imposed by friction and tight tolerances.
+The challenges of uncertainty become even more pronounced in contact-sensitive assembly tasks, such as those in traditional timber joinery, where success depends not only on geometric precision but also on controlling excessive force interactions arising from friction and tight tolerances.
 In such scenarios, even minor deviations introduced during fabrication or caused by material imperfections can easily lead to misalignment, resulting in jamming, stick-slip behavior, and material damage Apolinarska et al.
+[2021], Kramberger et al.
 [2022].
+2.2 Contact-rich manipulation of timber joints
 Handling force interactions in contact-rich robotic manipulation is often addressed through model-based impedance or admittance control laws Albu-Schäffer et al.
@@
 [2019], Schoettler et al.
-[2020] or hybrid approaches combining both Johannsmeier et al.
+[2020], or hybrid approaches combining both Johannsmeier et al.
 [2019], Hou et al.
 [2025].
-In the context of assembly with timber joinery, a few studies have explored learning-based methods; Apolinarska et al.
+In the context of assembly with timber joinery, a few studies have explored learning-based methods.
+Apolinarska et al.
 applied reinforcement learning for the motion control of an industrial robot to assemble single- and double-lap joints using robot pose and force/torque data from a wrist-mounted sensor Apolinarska et al.
 [2021].
-For the single-lap tasks, they also investigated the policy robustness by introducing angular and linear offsets to the initial robot pose, grasping the top piece.
+For the single-lap joints, they also investigated policy robustness by introducing angular and linear offsets to the initial robot pose while grasping the top piece.
 Their results showed strong performance in zero-offset cases and revealed challenges of policy adaptation when angular and linear offsets, as well as angled or double insertions, were introduced.
-The findings underscored the importance of robust policy learning for handling geometric variations and positional deviations in contact-rich assembly tasks.
+The findings underscored the importance of robust policy learning for handling geometric variations and positional deviations in contact-rich manipulation during assembly.
 Kramberger et al.
-also proposed a human–robot collaborative workflow using a learning-by-demonstration method that integrated a compliance controller into Cartesian-space dynamic movement primitives for perpendicular single-lap joint insertion; however, they relied on significantly loose tolerances Kramberger et al.
+also proposed a learning-by-demonstration method that integrated a compliance controller into Cartesian-space dynamic movement primitives for perpendicular single-lap joint insertion; however, they relied on significantly loose tolerances Kramberger et al.
 [2022].
 Other studies have explored semi-structured or open-loop methods for timber joinery assembly; Robeller et al.
-introduced a custom end-effector that generated vibration combined with a manually applied pulse force using a mallet to facilitate the connection of timber panels using dovetail joints Robeller et al.
+introduced a custom end-effector that generated vibration and combined it with a manually applied mallet-driven pulse force to facilitate the connection of wood panels using dovetail joints Robeller et al.
 [2017].
 Leung et al.
-demonstrated the use of an industrial robot in combination with manually placed distributed robotic clamps to assemble tight-fitting half-lap joints, highlighting the potential for precise and repeatable assembly.
-During their experiments, they observed challenges such as misalignments and collisions—largely arising from fabrication inaccuracies and deviations between digital and as-built models—that are common in construction-scale robotic applications Leung et al.
+demonstrated the use of an industrial robot, combined with manually placed distributed robotic clamps, to assemble tight-fitting half-lap joints, highlighting the potential for precise, repeatable assembly Leung et al.
 [2021].
+During their experiments, they observed challenges such as misalignments and collisions due to fabrication inaccuracies and deviations between digital and as-built models, which are common in construction-scale robotic applications.
 Rogeau employed visual feedback and fiducial markers for the assembly of wood panels with multiple mortise and tenon joints and reported successful full insertions in about 50% of their test cases across various joint configurations such as mortise tightness and tenon chamfer angles Rogeau [2023].
+2.3 Sensory–motor policy learning
 Model-based control can be particularly difficult when handling complex contact dynamics, such as those found in multi-body and multi-surface interactions or in the manipulation of deformable objects.
@@
 DDIMs extend this framework by formulating a non-Markovian reverse process, which allows the number of denoising steps used during inference to be reduced without changing the training procedure, thereby enabling significantly faster sampling.
-In the diffusion policy via robot action diffusion Chi et al.
+In diffusion policy via robot action diffusion Chi et al.
 [2024a], the denoising process generates a sequence of robot trajectories at each inference step.
-The approach is known to naturally capture multimodal action distributions, meaning it can represent multiple distinct yet equally valid action sequences for completing a task, as often occurs in human demonstrations.
+This approach is known to capture multimodal action distributions, meaning it can represent multiple distinct yet equally valid action sequences for completing a task, as often occurs in human demonstrations.
 The approach also maintains temporal coherence, meaning the policy predicts an entire sequence of future actions jointly rather than one step at a time.
-This joint prediction ensures that consecutive actions remain consistent (e.g., following one continuous path) rather than switching erratically between different predicted single actions, which can lead to unstable or jittery motion.
-Moreover, the method avoids the need for action discretization (breaking continuous movements into a fixed set of bins) or negative sampling (generating incorrect actions during training to guide learning) Chi et al.
+This joint prediction ensures that consecutive actions remain consistent rather than switching erratically between different predicted single actions, which can lead to unstable or jittery motion.
+Moreover, the method avoids the need for action discretization (breaking continuous movements into a fixed set of bins), or negative sampling (generating incorrect actions during training to guide learning) Chi et al.
 [2024a].
@@
 [2024a].
-Planning and control through contact remains a major challenge for both model- and learning-based approaches, due to the inherent complexity of contact dynamics Yang et al.
+2.4 Robot learning in construction
+Planning and control through contact remains a major challenge for both model- and learning-based approaches due to the inherent complexity of contact dynamics Yang et al.
 [2025].
-While diffusion policies have achieved notable success in various dexterous manipulation tasks, their effectiveness in contact-rich scenarios is still under active investigation Hou et al.
+While diffusion policies have achieved notable success in a range of dexterous manipulation tasks, their effectiveness in highly contact-rich scenarios is still under active investigation Hou et al.
 [2025], Kang et al.
 [2025], Wu et al.
+[2025], Zhou et al.
 [2026].
-Additionally, despite significant progress in construction robotic research, the application of robot learning for planning and control, particularly for contact-sensitive tasks, remains underexplored.
-In this study, we evaluate the performance and robustness of diffusion policies through a case study on the assembly of timber mortise and tenon joints under fabrication uncertainty.
-The following section outlines the objectives, novelty, and contributions of this research.
-1.2 Objectives and contributions
-In the research presented in this paper, we develop a sensory–motor diffusion policy learning framework via behavior cloning for the contact-rich robotic manipulation at construction scale.
-The study pursues three primary objectives: (1) Demonstrate applicability and performance of diffusion policies through a case study on the robotic assembly of mortise and tenon timber joints; (2) Investigate policy robustness under construction-relevant uncertainties, simulated as randomized perturbations to the mortise position; and (3) Identify actionable guidelines for data collection strategies and policy parameter tuning that yield the most successful and robust performance, providing a practical reference for future research and deployment.
-This paper presents one of the first systematic investigations in the AEC domain applying diffusion policy learning to robot planning and control.
-By enabling autonomous timber joinery assembly without reliance on hammering or human intervention, the proposed framework addresses physically demanding and craftsmanship-intensive tasks, offering a scalable learning framework for contact-sensitive manipulation tasks.
-The approach demonstrates potential for generalizing a wide spectrum of contact-rich assembly challenges, including complex timber joints, pipe fitting, and light metal framing.
-Ultimately, this work contributes to advancing robotic construction under uncertainty, bridging state-of-the-art generative policy learning with real-world AEC workflows, and provides both theoretical and practical advances that can accelerate adoption across the field.
+In particular, the deployment of these end-to-end sensory–motor policies on industrial manipulators for dexterous manipulation tasks has remained underexplored, likely due to compounded challenges arising from hardware constraints, high payloads, limited intrinsic compliance and force control, conservative safety-certified control interfaces, and restricted access to low-level control Kommey et al.
+[2025], Ruan et al.
+[2026], Delgado and Oyedele [2022].
+A growing body of work has investigated robot learning for construction tasks, most commonly through imitation learning and reinforcement learning Sun et al.
+[2026], Huang and Zou [2026], Duan et al.
+[2024, 2025], Yu et al.
+[2024], Li and Zou [2023], Apolinarska et al.
+[2021].
+Notably, most of these studies were conducted on lightweight collaborative or mobile manipulators operating under relatively compliant hardware conditions.
+Among these, Sun et al.
+investigated rebar slot insertion and rebar tying using a mobile manipulator, combining visual servoing for coarse positioning with imitation learning for precise execution Sun et al.
+[2026].
+The rebar insertion task required precise alignment of the stirrup with three slots prior to insertion, with a reported total insertion clearance of approximately 1.5 mm.
+Similarly, Duan et al.
+studied a cable-in-duct installation task using a zero-shot Sim2Real visual–tactile reinforcement learning framework, evaluating robustness by varying pipe diameters (16–25 mm) for a 4 mm fish tape Duan et al.
+[2025].
+Among learning-based studies in construction, Apolinarska et al.
+(mentioned in Section 2.2) is one of the few studies deploying reinforcement learning on a large-scale industrial manipulator, evaluating millimeter-scale (1 and 2 mm) geometric clearance for timber lap joint assembly Apolinarska et al.
+[2021].
+Compared to these prior works, the present study investigates sensory–motor policy learning on industrial manipulators for timber joinery assembly under substantially tighter sub-millimeter clearances and positional perturbations that exceed these clearances, thereby probing the limits of policy robustness in a structurally constrained, contact-rich manipulation task.
 3 Methods
-This section outlines the experimental setup, data collection process, policy learning approach, and evaluation procedure used in our robotic manipulation system for timber joinery.
-We implement diffusion policy learning via behavioral cloning using robot pose and F/T data collected from teleoperated demonstrations.
+This section outlines the experimental setup, including the multi-robot workcell, communication stack, and teleoperation system to collect expert demonstrations (Section 3.1), data synchronization and processing methods to prepare the dataset for training (Section 3.2), a detailed overview of the sensory–motor diffusion policy training method (Section 3.3), and the policy inference details to evaluate the trained policies on our experimental setup (Section 3.4).
 Figure 2: Multi-robot setup consisting of two 6-axis industrial robotic arms.
@@
 Our experimental setup consists of two primary systems for conducting this research.
-The first is a multi-robot setup equipped with customized end effectors for executing the timber assembly task.
-The second is a teleoperation system for manually operating the robotic setup and collecting data for demonstrations.
+The first is a multi-robot setup equipped with customized end effectors for contact-rich manipulation and timber assembly.
+The second is a teleoperation system for manually operating the robotic setup and collecting data during demonstrations.
 This subsection details the specific hardware used for our experiments and their software interfaces.
-The multi-robot setup consists of two 6-axis industrial robotic arms [^1] ABB IRB 4600 ABB Group , each with a 40 kg payload and a 2.55 m reach mounted on linear tracks (Fig.
+The multi-robot setup consists of two six-axis industrial robotic arms [^1] ABB IRB 4600 ABB Group , each with a 40 kg payload and a 2.55 m reach mounted on linear tracks (Fig.
 2).
-Both robotic arms utilize custom pneumatically-driven end effectors designed for gripping long timber elements with varying cross-sectional profiles, as typically encountered in timber frame construction.
-One arm grasps the timber mortise in place and can systematically simulate construction deviations by adding positional offsets.
-The other arm grasps the timber tenon and is the primary agent for manipulation and completing the assembly task.
-Because the industrial robotic arms do not provide native access to joint torque values, we developed the end effector on the tenon-grasping arm to include a 6-axis F/T sensor [^2] ATI Delta IP60, with SI-330-30 calibration ATI Industrial Automation to enable F/T data collection (Fig.
+Both robotic arms utilize custom pneumatically-driven end effectors designed for gripping long elements with varying cross-sectional profiles.
+Because the industrial robotic arms do not provide native access to joint torque values, one of the end effectors includes a six-axis F/T sensor [^2] ATI Delta IP60, with SI-330-30 calibration ATI Industrial Automation to enable F/T data streaming (Fig.
 3).
-For safety, the end effector is also equipped with a pneumatic anti-collision sensor [^3] Schunk OPR 081-P00 Schunk , which both provides passive mechanical compliance and triggers a protective stop on the robotic arm under excessive force, safeguarding the robotic setup and F/T sensor.
+For safety, the same end effector is also equipped with a pneumatic anti-collision sensor [^3] Schunk OPR 081-P00 Schunk , which both provides passive mechanical compliance and triggers a protective stop on the robotic arm under excessive force, safeguarding the robotic setup and F/T sensor.
 Motion control and pose feedback for the robotic arms utilize ABB’s Externally Guided Motion (EGM) feature ABB Robotics [2017], a UDP-based real-time control interface.
 F/T data is transferred over EtherCAT to a central Programmable Logic Controller (PLC) [^4] Beckhoff CX2062 Beckoff .
-All data communications are centralized on a desktop computer employing ROS [^5] ROS 2 Jazzy Jalisco Open Robotics [2023] interfaces and Python wrappers.
-Figure 3: Gripper for holding the tenon piece equipped with anti-collision and F/T sensors.
+All data communications are centralized on a desktop computer connected to the central PLC via a local area network (LAN) employing ROS [^5] ROS 2 Jazzy Jalisco Open Robotics [2023] interfaces and Python wrappers.
+Figure 3: Custom gripper end effector equipped with anti-collision and F/T sensors.
 The teleoperation system (Fig.
-4) consists of a virtual reality (VR) system [^6] HTC VIVE Pro 2 HTC Corporation , and we developed a unilateral pose-to-pose transformation from the VR controller to the robotic arm’s end effector.
-This transformation is essential for aligning the coordinate frames of the VR system and the robotic setup, and for compensating for the human demonstrator’s orientation relative to the robot.
-For instance, when the demonstrator is facing the robotic arm, pushing the controller forward should consistently move the end effector away from the demonstrator, regardless of their position.
-In addition, this transformation includes an adjustable scaling parameter that maps the demonstrator’s control motions to proportionally smaller movements on the robot, enhancing the demonstrator’s fine motor control capabilities during high-precision tasks such as timber joint assembly.
-During teleoperation, robot motion and data collection (see Section 2.2) are enabled by the hand controller’s trigger.
-Similar to clicking and dragging a mouse, holding the trigger moves the robot in the desired direction, and releasing it allows the controller to be repositioned without affecting the robot or recording idle data.
-This allows for long motions to be performed in smaller, controlled segments.
-The VR controller interface utilizes OpenVR Valve Corporation [2024] in Python and ROS, integrating directly with the multi-robot setup by streaming mapped poses to EGM for fluid, real-time teleoperation.
+4) implemented in this study is based on a virtual reality (VR) interface and follows the framework previously introduced in our prior work Ruan et al.
+[2026].
+In this system, six-degree-of-freedom hand poses are streamed from the VR hand controller [^6] HTC VIVE Pro 2 HTC Corporation to the desktop computer via the OpenVR Valve Corporation [2024] application programming interface (API) and mapped in real time to robot end-effector commands streamed over EGM.
+Specifically, we implement a unilateral pose-to-pose transformation that aligns the VR controller frame with the robot tool center point (TCP) frame.
+This transformation ensures consistent directional behavior between human input and robot motion (e.g., forward controller motion always results in motion away from the operator, regardless of physical orientation).
+When the operator presses the controller trigger, the current hand pose and robot TCP pose are recorded as reference frames.
+While the trigger remains engaged, subsequent controller motions are interpreted as relative displacements with respect to these references and converted into incremental robot motion commands.
+Translational and rotational components are processed separately and may be independently scaled, enabling fine-grained control during high-precision, contact-rich operations such as tight-tolerance insertion.
+Additionally, the operator receives haptic feedback through controller vibrations when the measured force from the F/T sensor exceeds a predefined threshold (less than the sensing range maximum value).
+This cue provides awareness of contact intensity during teleoperation, supporting safer interaction and more controlled force application.
 Figure 4: VR-based teleoperation system for collecting human expert demonstrations.
-2.2 Data collection
-To train our diffusion policies, we created datasets consisting of expert demonstrations.
-An overview of the data collection workflow is illustrated in Fig.
-5.
-During data collection, a human demonstrator utilized the teleoperation system to manipulate a vertical tenon into the mortise.
-The demonstrator initiates the task by simultaneously straightening the tenon and executing a downward motion to guide it into the mortise.
-To improve the robustness of the learned policy under fabrication uncertainties, we dedicated approximately 25% of the demonstrations to error recovery scenarios.
-In these cases, the demonstrator intentionally initiates contact with the edges of the mortise and performs a sliding maneuver along the surface to realign and insert the tenon, mimicking realistic recovery behaviors in the presence of misalignment.
-Each collected demonstration consists of timestamped robot poses and F/T data during motion (while the VR controller’s trigger is held).
-Figure 5: Data collection and policy rollout workflow.
-The raw data for each demonstration is recorded in real-time for each sensor interface: pose data is collected from EGM as a seven-dimensional (7D) vector (3D position in millimeters + 4D quaternion rotation of the tenon gripper end effector) every 12 ms (approximately 83 Hz), while F/T data is collected as a 6D vector (3D force in newtons + 3D torque in newton-meters) at 64 Hz after passing through an infinite impulse response (IIR) low-pass filter to attenuate frequency components above 64 Hz.
+3.2 Data processing
+Figure 5: Data collection via teleoperation.
+During teleoperation, the raw data for each demonstration is recorded for each sensor interface (Fig.
+5): pose data is collected from the robot controller as a seven-dimensional (7D) vector (3D position in millimeters + 4D quaternion rotation of the tenon gripper end effector) every 12 ms (approximately 83 Hz), while F/T data is collected as a 6D vector (3D force in newtons + 3D torque in newton-meters) at 64 Hz after passing through an infinite impulse response (IIR) low-pass filter to attenuate frequency components above 64 Hz.
 This filtering step mitigates aliasing and improves signal quality by removing sensor noise and dynamic vibrations introduced by high-speed motion or structural resonances.
-To temporally align the different sample rates to a common time base, the pose and F/T data are interpolated to 60 Hz; position and F/T data are linearly interpolated, while rotation data uses spherical linear interpolation to preserve smooth interpolation.
-Finally, each episode is post-processed to reduce unwanted noise in the demonstrations and remove periods of inactivity or unintentional stillness.
+To temporally align the different sample rates to a common time grid, the pose and F/T data are interpolated to 60 Hz; position and F/T data are linearly interpolated, while rotation data uses spherical linear interpolation to preserve smooth interpolation.
+Finally, each episode is post-processed to remove periods of inactivity or unintentional stillness and reduce unwanted noise in the demonstrations.
 Removing idle or near-idle periods is critical as it helps the policy training to focus on active, meaningful interactions rather than noise or moments where no significant movement occurs.
@@
 In contrast, we used a 1st-order low-pass Butterworth filter with a higher 10 Hz cutoff to preserve sharper transitions (e.g., at the moment of contact) while effectively removing high-frequency noise that may come from mechanical oscillations or sensor inaccuracies.
-Actions are then derived from finite differences between successive pose data points, and any resulting vectors with magnitudes below a set threshold are removed to discard idle or near-idle actions.
 Figure 6: Data filtering using a low-pass Butterworth filter.
@@
 Only the first 3 dimensions for each pose and F/T data are shown.
-2.3 Policy learning
-Our research employs CNN-based diffusion policy learning via behavioral cloning Chi et al.
+3.3 Policy training
+We train a CNN-based sensory–motor diffusion policy via action diffusion Chi et al.
 [2024a].
-The method is based on DDPMs Sohl-Dickstein et al.
-[2015], Ho et al.
-[2020] and DDIMs Song et al.
-[2022], using DDPM for the forward noising process during training and DDIM for the reverse denoising process and faster inference during rollouts.
-Our diffusion policies predict a sequence of robot actions (i.e., the robot pose) conditional on the observation horizon of robot pose and F/T data.
-The foundation for our diffusion models is built upon the work of Bogert et al.
+The policy training uses a standard DDPM-style training Ho et al.
+[2020].
+During inference, we adopt a DDIM-style sampling procedure Song et al.
+[2022], which enables accelerated trajectory generation using fewer denoising steps for efficient action prediction during robot execution.
+The policy predicts a sequence of robot actions (i.e., the robot TCP pose) conditional on an observation horizon of robot pose and F/T data.
+The foundation for our diffusion policy is built upon the work of Bogert et al.
 van den Bogert et al.
 [2024].
-When loading the dataset for training our diffusion policies, trajectories of 64 points are subsampled from each demonstration.
+When loading the dataset for training the diffusion policies, trajectories of 64 points are subsampled from each demonstration.
 This subsampling normalizes the trajectory length for each demonstration, as well as enables faster training iterations.
-Next, we converted pose quaternions to a continuous 6D rotation representation (by first converting to rotation matrices then concatenating the first two columns) to improve training stability Zhou et al.
+We also convert pose quaternions to a continuous 6D rotation representation (by first converting to rotation matrices, then concatenating the first two columns) to improve training stability Zhou et al.
 [2019], Chi et al.
 [2024a], resulting in 9D pose vectors.
-We then applied min-max normalization to scale all features to the range [-1, 1].
+We then apply min-max normalization to scale all features to the range [-1, 1].
+Below, we formalize the diffusion policy training on the subsampled and normalized dataset.
 Given a pose observation horizon $T_{o}^{p}$, F/T observation horizon $T_{o}^{f}$, and action prediction horizon $T_{p}$, each training sample at timestep $t$ consists of an observation feature $\mathbf{O}_{t}$ concatenating the last $T_{o}^{p}$ and $T_{o}^{f}$ and a corresponding action sequence $\mathbf{A}_{t}$ consisting of the next $T_{p}$ absolute poses.
-During the rollouts, $\mathbf{A}_{t}$ is predicted conditioned on $\mathbf{O}_{t}$.
+During inference, $\mathbf{A}_{t}$ is predicted conditioned on $\mathbf{O}_{t}$.
 The reverse denoising process for action prediction starts from ${A}^{K}$ sampled from Gaussian noise.
@@
 As implemented in Chi et al.
-[2024a], we add the sampled random noise $\epsilon^{k}$ to the noise-free actions $\mathbf{A}_{0}$ in $K$ forward steps (original DDPM method).
+[2024a], the sampled random noise $\epsilon^{k}$ is added to the noise-free actions $\mathbf{A}_{0}$ in $K$ forward steps (original DDPM method).
 Policy training was conducted on high-performance computing (HPC) clusters [^7] Princeton Research Computing’s Della cluster, which provide powerful computational resources for efficiently processing large datasets and enabling faster, scalable learning by leveraging parallel computing capabilities.
-Each policy was trained for a fixed amount of time (1 hour), with each training lasting at least 1,000 epochs.
+Each policy was trained for a fixed amount of time (1 hour), with each training using a batch size of 128 and lasting at least 1,000 epochs.
 Only the model weights corresponding to the epoch with the lowest validation loss were used to evaluate performance during rollouts.
-To account for training variance, we trained 4 models for each set of parameters and averaged their success rates when computing their performance (see Section 2.4).
-2.4 Evaluation
+3.4 Policy inference
 To evaluate the performance of our trained policies, we perform rollouts on the experimental setup (Section 3.1).
-The starting position for the tenon piece during all rollouts is fixed and identical to the demonstrations.
-During policy rollouts, we specify additional parameters for the number of inference steps $K_{\text{inf}}$, DDIM/DDPM interpolation $\eta$ Song et al.
+During policy inference, we specify additional parameters for the number of inference steps $K_{\text{inf}}$, DDIM/DDPM interpolation $\eta$ Song et al.
 [2022], and action execution steps $T_{a}$.
@@
 The amount of stochasticity during the reverse denoising process is controlled by the interpolation parameter $\eta\in[0,1]$, where $\eta=0$ is the DDIM process and $\eta=1$ is the original DDPM process.
-This means that when $\eta=0$, the standard deviation of the noise added during sampling is 0, meaning that the reverse process is fully deterministic, producing high-fidelity imitations of demonstrations, and when $\eta=1$, the reverse process is fully stochastic, introducing broader behavioral coverage with exploration and variability.
+This means that when $\eta=0$, the standard deviation of the noise added during sampling is 0, making the reverse process fully deterministic and producing high-fidelity imitations of demonstrations.
+When $\eta=1$, the reverse process is fully stochastic, introducing broader behavioral coverage through exploration and variability.
 The number of action execution steps $T_{a}\leq T_{p}$ denotes that of the $T_{p}$ predicted actions at each time step; we execute the first $T_{a}$ actions.
-Decreasing the value of $T_{a}$ increases the frequency of replanning over a receding horizon, increasing the responsiveness of the system; however, it can also potentially decrease the speed of rollouts by spending more time on inference.
-For values of $T_{a}>2$, we apply an average pooling with a kernel size of $T_{a}-1$ across each dimension of the actions, smoothing the action sequence.
-In contrast to data collection during demonstrations, observations during rollouts are not temporally aligned and instead utilize the latest available pose and F/T data when the observation is retrieved.
+Decreasing the value of $T_{a}$ increases the frequency of replanning over the receding horizon, thereby increasing the system’s responsiveness; however, it may also reduce the speed of rollouts by increasing inference time.
+For values of $T_{a}>2$, we apply average pooling with kernel size $T_{a}-1$ across each action dimension, thereby smoothing the action sequence.
+In contrast to data collection during demonstrations, observations during inference are not temporally aligned, and instead, we use the latest available pose and F/T data when the observation is retrieved.
 These observations are normalized using the same normalization parameters as the dataset before being passed to the policy for inference.
 The output action sequence is similarly denormalized back into the original scale using the same action normalization parameters during training.
-These predicted actions are then sent through EGM to the robot controller, and we use a similar communication protocol to the data collection to receive pose and F/T feedback (see Fig.
-5).
-The policy performance is determined by rolling out 5 times for each of the 4 trained models with fixed parameters, to account for stochasticity during test time, resulting in an average success rate (Avg.
+4 Experiments
+This section describes the experimental methodology and evaluation protocol used in this study.
+We first introduce the mortise and tenon assembly task and the criteria used to measure policy performance.
+We then present the data collection procedure used to construct the demonstration datasets.
+Finally, we outline two experimental phases: Phase 1 focuses on baseline evaluation and parameter initialization under fixed conditions, while Phase 2 fine-tunes the parameters, examines policy robustness under randomized mortise position perturbations, including additional force ablation and demonstration count studies.
+4.1 Task Description
+The experimental task is the contact-rich robotic assembly of a vertical mortise and tenon joint as illustrated in Fig.
+7.
+Each robot initially grasps a timber element, and we assume that a conventional motion planner is responsible for placing the grasped tenon at a pre-insertion pose.
+This pre-insertion pose is constant for all demonstrations and rollouts, located near the mortise, positioned approximately 15 mm above the mortise hole, and subject to an initial angular misalignment of approximately 6 degrees about the y-axis (Fig.
+7).
+Uncertainties associated with grasping are outside the scope of this study and are not considered.
+The task objective is to fully insert the tenon into the mortise using compliant, contact-aware motion.
+Only the robot with the grasped tenon is controlled while the mortise pose remains fixed throughout the task.
+The geometric dimensions of the mortise and tenon used in all experiments are shown in Fig.
+7.
+The clearance between the tenon cheeks and the mortise walls is 0.1 mm, representing a tight structural fit typical of traditional timber joinery, while the clearance between the tenon edges and the mortise walls is approximately 2 mm.
+The tight tolerance of the tenon cheeks governs load transfer and frictional contact during insertion and is the primary source of contact sensitivity in the task.
+Figure 7 also visualizes the boundary corresponding to the 5 mm and 10 mm offsets used in the Phase 2 experiments (Section 4.4).
+These offsets mimic realistic deviations arising from positioning errors, tolerance accumulation, and material imperfections commonly encountered in construction-scale assembly.
+Figure 7: Mortise and tenon assembly task (left), and terminology, dimensions, and reference frame (right).
+For each experimental configuration, the policy is trained 4 times with randomized seeds, and policy performance is determined by rolling out 5 times for each trained model, to account for stochasticity during test time, yielding an average success rate (Avg.
 SR) across 20 combined rollouts.
 A rollout is recorded as a success if the tenon element fully inserts into the mortise hole without triggering any collision errors (i.e., excessive force at either the collision sensor or the robot joints).
-A rollout is recorded as a failure if any collision errors are triggered by the industrial robot controller or if the policy does not appear to be making any meaningful progress towards task completion, as determined by the operator.
-3 Experiments
-We conducted a two-phase study: in Phase 1 we performed static experiments with a fixed mortise location to establish a general performance evaluation for the diffusion policy in our industrial setup; in Phase 2 we performed uncertainty experiments with randomized hole offsets to evaluate the robustness of the diffusion policies.
-The dimensions of the mortise and tenon we used for the experiments are illustrated in Fig.
-7.
-The tolerance between the tenon cheeks and the mortise walls was 0.1 mm, and the tolerance between the tenon edges and the mortise walls was about 2 mm.
-The tight tolerance of the tenon cheeks is what ensures a structural fit for a mortise and tenon joint.
-This figure also visualizes the boundary of the 5 mm and 10 mm offsets used in the uncertainty experiments (see Section 3.2).
-The tenon is aligned along the world’s z-axis, while the mortise piece lies horizontally along the x-axis.
-Each demonstration and rollout begins with the tenon element grasped at a fixed starting pose, positioned approximately 15 mm vertically above the mortise hole with an approximately 6°angular deviation about the y-axis.
-Figure 7: Mortise and tenon terminology, dimensions, and reference frame.
+A rollout is recorded as a failure if the industrial robot controller triggers any collision errors or the policy does not appear to make meaningful progress toward task completion, as determined by the operator (see examples of successful and failed rollouts in Fig.
+9).
+4.2 Data collection
+To train the diffusion policies, we created two datasets consisting of expert demonstrations.
+During data collection, a human demonstrator utilized the teleoperation system (Section 3.1) to complete the vertical mortise and tenon assembly task.
+The first dataset consists of 100 demonstrations with fixed starting positions for the mortise and tenon.
+This dataset is used by the Phase 1 experiments to initialize parameter tuning (Section 4.3).
+The second dataset consists of 400 demonstrations with a fixed tenon starting position but a randomized mortise position, simulating fabrication uncertainty.
+The mortise position was perturbed by applying a planar translational offset sampled in polar coordinates, with direction $\theta\sim U(0,2\pi)$ and magnitude $r\sim U(0,10\text{~mm})$, and converted to Cartesian offsets $\left(\Delta x,\Delta y\right)=(r\cos{\theta},r\sin{\theta})$.
+This sampling scheme was selected to approximate assembly tolerances, placing greater probability mass near small offsets while still demonstrating larger deviations and boundary cases, and remaining simple to implement directly on the industrial robot controller.
+To improve the robustness of the learned policy under fabrication uncertainties, we dedicated 100 (25%) of the demonstrations to error recovery scenarios Black et al.
+[2026].
+In these cases, the demonstrator intentionally initiates contact with the edges of the mortise and slides along the surface to realign and insert the tenon, mimicking realistic recovery behaviors in the presence of misalignment.
+This dataset is used in the Phase 2 experiments (Section 4.4).
 4.3 Phase 1: Static experiments
-The objective of the Phase 1 experiments is to show the application and performance of diffusion policies in a building-scale robotic assembly setting.
-For this experiment, we created a dataset of 100 demonstrations following the data collection method described in Section 2.2, with the exception that no demonstrations included error recovery cases and demonstrations were not post-processed.
-To simplify the assembly task, the mortise position was held fixed across all demonstrations and during rollouts.
-Through repeated experimentation, we identified the best-performing set of parameters (reported in Section 4.1), which were then used to initialize parameter tuning in the Phase 2 experiments described in the following subsection.
+The objective of the Phase 1 experiments is to evaluate the applicability and baseline performance of diffusion policies for contact-rich manipulation of the mortise and tenon joint.
+The policies were trained on the first dataset of 100 demonstrations (Section 4.2).
+Through repeated experimentation, we identified the best-performing parameter set (reported in Section 5.1) and used it to initialize parameter tuning in the Phase 2 experiments described in the following subsection.
 4.4 Phase 2: Uncertainty experiments
-The objective of the Phase 2 experiments is to demonstrate the robustness of diffusion policies in handling fabrication uncertainties for contact-rich timber joinery.
-More specifically, we evaluate how the policies perform when the mortise is subjected to random position offsets simulating fabrication tolerances.
-We created a dataset of 400 demonstrations following the data collection method outlined in Section 2.2.
-To simulate fabrication tolerances, the position of the mortise in these demonstrations was perturbed by a displacement vector sampled in polar coordinates.
-The angular component was drawn uniformly from $0,2\pi)$ radians, and the radial component was drawn uniformly from $[0,10]$ mm.
-Unlike uniform sampling over the area of a disk, this approach yields a higher density of samples near the origin, biasing the distribution toward smaller offsets while still allowing for the possibility of larger deviations.
-During rollouts, we evaluated each policy at three goal offset distances: 0 mm (no offset), 5 mm, and 10 mm.
+The objective of the Phase 2 experiments is to evaluate the robustness of diffusion policies in handling uncertainties arising from positioning errors, tolerance accumulation, and material imperfections commonly encountered in construction-scale assembly.
+Specifically, we assess policy performance when the mortise is subjected to randomized positional offsets to simulate fabrication uncertainty.
+The mortise orientation is not randomized and remains constant.
+The policies for these experiments were trained using the second dataset of 400 demonstrations (Section 4.2).
+During rollouts, we evaluated each policy with varying mortise position perturbations: 0 mm (no perturbation), 5 mm, and 10 mm.
 These values were selected to evaluate the policies under no uncertainty (0 mm), uncertainty within the demonstration distribution (5 mm), and uncertainty at the edge of the demonstration distribution (10 mm).
-For each offset condition, the goal position was randomized uniformly along the circumference of a circle with a radius of the specified distance.
-The success rate of a policy was computed separately for each offset distance, and its overall performance was reported as the average success rate across all three conditions.
-In the first part of the Phase 2 experiments, we systematically tuned the parameters of the diffusion policy to maximize the success rate.
+For each offset condition, the mortise position was uniformly randomized along the circumference of a circle with a radius equal to the specified distance.
+The Avg.
+SR of a policy was computed separately for each offset distance, and its overall performance was reported as the average total success rate (Avg.
+total SR) across all three conditions.
+In the first part of the Phase 2 experiments, we systematically tuned the diffusion policy parameters to maximize the success rate.
 Starting from a candidate configuration derived from the Phase 1 experiments (Section 4.3), we performed sequential, greedy parameter tuning, optimizing one parameter at a time while holding others fixed.
@@
 The impact of individual parameters on policy performance is reported in Section 5.2.1.
-In the second part of the Phase 2 experiments, we conducted two ablation studies to assess the impact of F/T observations and demonstration count on policy performance.
-In the F/T ablation study, we evaluated the performance of the full model with F/T inputs masked (i.e., input F/T values are set to 0), as well as a variant trained using only pose observations.
-For the demonstration ablation, we trained policies using 25, 50, 100, and 200 demonstrations, while keeping all other parameters identical to the full model.
-The results of these studies are presented in Sections 4.2.2 and 4.2.3.
+In the second part of the Phase 2 experiments, we conducted two studies to assess the impact of F/T observations and demonstration count on policy performance.
+In the F/T ablation study, we evaluated the performance of the full model with F/T inputs masked (i.e., F/T values set to 0) during inference, as well as a variant trained only on pose observations.
+For the demonstration count, we trained policies using 25, 50, 100, and 200 demonstrations (each including 25% error recovery cases as explained in Section 4.2), while keeping all other parameters identical to the full model.
+The results of these studies are presented in Sections 5.2.2 and 5.2.3.
 5 Results and discussion
@@
 The loss curves were first averaged across all 4 training iterations, then smoothed using a Gaussian filter with a standard deviation $\sigma=10$ for visualization.
-Figure 8: Smoothed mean squared error (MSE) training and validation loss curves for Policy 4, averaged then smoothed across all training iterations.
+Figure 8: Smoothed mean squared error (MSE) training and validation loss curves for Policy 4, averaged and then smoothed across all training iterations.
 The shaded area represents $\pm 1$ standard deviation.
@@
 5.1 Phase 1: Static experiments
-After repeated testing and parameter tuning, the policy achieved a 100% average success rate over 20 rollouts using the parameter values listed in Table 1.
+In Phase 1, after repeated testing and initial parameter tuning, the policy achieved a 100% average success rate over 20 rollouts using the parameter values listed in Table 1.
 This performance demonstrates the capability of diffusion policies to reliably assemble a timber mortise and tenon joint under no uncertainty.
@@
 $\eta$ DDIM/DDPM interpolation 0.5
-During the first phase of the experiments, we also evaluated encoder networks for pose and F/T inputs to the diffusion policy.
+During the Phase 1 experiments, we also evaluated encoder networks for pose and F/T inputs to the diffusion policy.
 Each encoder consisted of a single fully connected layer that projected the raw input vector to a latent representation of fixed length (e.g., 64 or 128 units).
@@
 Future research should explicitly account for these fabrication-induced variations when developing and evaluating assembly policies.
-4.2 Phase 2: Uncertainty experiments
-4.2.1 Full model parameter tuning
-The sequential parameter tuning in Phase 2 was limited to three stages: (1) pose and F/T observation horizons; (2) action prediction horizon; (3) number of inference steps and DDIM/DDPM interpolation.
-The learning rate, weight decay, and forward noising steps were fixed at the values determined in Phase 1.
-Table 2 summarizes the results from the first two stages.
-The highest average total success rate of 75%, computed over the three mortise offset conditions, was achieved using an observation horizon of 1 for both pose and F/T modalities combined with a longer action prediction horizon of 16.
-This configuration suggests that, for the uncertainty conditions tested, the policy benefited from focusing on the most recent state information while planning actions further into the future, potentially allowing it to better anticipate and correct for deviations introduced by positional offsets.
 Table 2: Success rates for training parameter tuning.
@@
   • D = randomized hole offset.
-Table 3 summarizes the final stage of rollout parameter tuning, evaluating different combinations of inference steps and DDIM/DDPM interpolation values.
-These tests were conducted only for mortise offsets of 5 mm.
-The initial parameter values selected from the Phase 1 experiments achieved the highest performance, with a 65% average success rate, substantially outperforming other tested configurations.
-Table 3: Success rates for rollout parameter tuning.
+5.2 Phase 2: Uncertainty experiments
+In Phase 2, we first perform systematic, sequential parameter tuning, starting from the Phase 1 configuration to identify the best-performing diffusion policy (referred to as the full model).
+Second, we analyze the impact of F/T observations and the number of demonstrations on policy performance.
+5.2.1 Full model parameter tuning
+The sequential parameter tuning during the Phase 2 experiments was limited to three stages: 1) pose and F/T observation horizons; 2) action prediction horizon; 3) number of inference steps and DDIM/DDPM interpolation.
+The learning rate, weight decay, and forward noising steps were fixed at the values determined during the Phase 1 experiments.
+Table 2 summarizes the results from the first two stages.
+The highest average total success rate of 75%, computed over the three mortise offset conditions, was achieved using an observation horizon of 1 for both pose and F/T modalities combined with a longer action prediction horizon of 16.
+This configuration suggests that, for the uncertainty conditions tested, the policy benefited from focusing on the most recent state information while planning actions further into the future, potentially allowing it to better anticipate and correct for deviations introduced by positional offsets.
+Table 3 summarizes the final stage of parameter tuning, evaluating different combinations of inference step $K_{inf}$ and DDIM/DDPM interpolation $\eta$.
+These tests were conducted only for mortise offsets of 5 mm, as the overall trend in success rates was consistent across other offsets.
+The initial parameter values selected from the Phase 1 experiments (i.e., $K_{inf}$ = 32 and $\eta$ = 0.5) achieved the highest performance, with a 65% average success rate, substantially outperforming other tested configurations.
+Table 3: Success rates for inference parameter tuning.
 Policy # Demos $T_{p}$ $T^{p}_{o}$ $T^{f}_{o}$ $T_{a}$ D (mm) $K_{\text{inf}}$ $\eta$ Avg. SR (%)
@@
   • D = randomized hole offset.
-4.2.2 Force/torque ablation study
+5.2.2 Force/torque ablation
 Table 4 summarizes the results of the F/T ablation study, and Fig.
@@
 10) for this condition suggests that these successes were likely the result of random factors rather than reliable inference.
-This highlights the critical role of force feedback in consistent policy performance, particularly under larger fabrication uncertainties.
+This highlights the critical role of force feedback in consistent policy performance, particularly under larger perturbations.
 Table 4: Success rates for the F/T ablation study.
@@
   • D = randomized hole offset; PF = models trained with pose and force data; P = models trained with pose-only data.
-  • *The forces were masked during rollouts.
+  • *The forces were masked during inference.
 Figure 10: Success rates of diffusion policies with force masked (PF*) and trained only on pose data (P) compared to the full model (PF), evaluated at three mortise offsets (0, 5, and 10 mm).
 Error bars indicate the standard error of the mean (SEM) computed across the 4 independently trained models for each set of parameters.
-4.2.3 Demonstration quantity ablation study
-Table 5 summarizes the results of the demonstration quantity ablation study.
+5.2.3 Demonstration count
+Table 5 summarizes the results of the demonstration quantity study.
 Policies trained on only 25 demonstrations failed to fully converge and achieved no success during rollouts, and were therefore excluded from the reported results.
@@
 Notably, the total number of demonstrations required to achieve robust performance was much higher than anticipated.
-This is likely attributable to the trajectory subsampling strategy used during policy learning (Section 2.3), which significantly reduced the total number of training samples available to the policy.
-Future research should utilize alternative data preprocessing strategies to maintain a sufficient quantity of training data to enable reliable policy learning.
-Table 5: Success rates for the demonstration quantity ablation study.
+This is likely attributable to the trajectory subsampling strategy used during policy training (Section 3.3), which significantly reduced the total number of training samples available to the policy.
+Future research should utilize alternative data preprocessing strategies to maintain a sufficient quantity of training data to enable more reliable policy learning.
+Table 5: Success rates for various demonstration counts.
 Policy # Demos $T_{p}$ $T^{p}_{o}$ $T^{f}_{o}$ $K_{\text{inf}}$ $\eta$ $T_{a}$ D (mm) Avg. SR (%) Avg. total SR (%)
@@
 6 Conclusion
-The experimental results presented in this work demonstrate that diffusion policies can achieve stable convergence and high success rates in a building-scale contact-rich assembly task representative of real-world construction.
-In Phase 1, we established a baseline by identifying parameter settings capable of achieving 100% average success rate in a deterministic mortise and tenon joint assembly task.
-In Phase 2, we evaluated policy robustness under controlled fabrication uncertainties, identifying parameter configurations that improved generalization to offsets of the mortise position, quantifying the critical role of force/torque feedback, and determining the demonstration quantity threshold necessary for reliable performance.
-Together, these findings address the study’s core objectives: they validate the applicability of diffusion policies to building-scale robotic assembly and systematically characterize their robustness under fabrication uncertainty.
+The experimental results presented in this work demonstrate that diffusion policies can achieve stable convergence and high success rates in a contact-rich assembly task representative of real-world construction using industrial robotic arms.
+In the Phase 1 experiments, we established a baseline by identifying parameter settings capable of achieving 100% average success rate in a deterministic mortise and tenon joint assembly task.
+In the Phase 2 experiments, we evaluated policy robustness under controlled mortise position perturbations, identifying parameter configurations that improved generalization to offsets of the mortise position, quantifying the critical role of force/torque feedback, and determining the demonstration quantity threshold necessary for reliable performance.
+These findings validate the applicability of diffusion policies to building-scale robotic assembly using industrial robots and systematically characterize their robustness under fabrication uncertainty.
 Beyond the performative metrics, the experiments also yielded actionable insights for parameter tuning, sensor fusion, and demonstration count, forming a generalizable methodology for future contact-rich construction and assembly tasks.
-By demonstrating that diffusion policies can be systemically adapted and tuned to handle fabrication uncertainties without manual intervention, this research establishes a scalable, feedback-driven robot learning framework that addresses not only the technical demands of precision assembly, but also broader industry challenges such as labor shortages and the need to reduce reliance on physically taxing, repetitive, and high-skill manual work.
+By demonstrating that diffusion policies can be systemically adapted and tuned to handle fabrication uncertainties without manual intervention, this research establishes a scalable, feedback-driven robot learning framework that addresses not only the technical demands of precision assembly but also broader industry challenges such as labor shortages and the need to reduce reliance on physically taxing, repetitive, and high-skill manual work.
 Beyond timber joinery, the presented methodology can be generalized to a wide range of construction and manufacturing scenarios, from complex timber connections and pipe fitting to light metal framing, where expert craftsmanship has traditionally been essential.
@@
 6.1 Limitations and future work
-While this work demonstrates the feasibility and robustness of applying sensory–motor diffusion policy learning to contact-rich robotic assembly at construction scale, several limitations remain; our approach relies exclusively on robot pose and force data to train the policy and evaluate robustness, whereas diffusion policies for robot planning and control are commonly formulated as visuomotor policies using end-to-end mappings from RGB or tactile images to robot commands Chi et al.
+While this work demonstrated the feasibility and robustness of applying sensory–motor diffusion policy learning to contact-rich robotic assembly at construction scale, several limitations remain; our approach relies exclusively on robot pose and force data to train the policy and evaluate robustness, whereas diffusion policies for robot planning and control are commonly formulated as visuomotor policies using end-to-end mappings from RGB or tactile images to robot commands Chi et al.
 [2024a, b].
-Our simplification of the sensory input space increases training speed but may limit the model’s ability to generalize to visually complex or occlusion-prone construction scenarios, such as joint assembly in cluttered framing environments or alignment tasks in multi-material panel installations.
-Moreover, incorporating expressive state-of-the-art learning models, such as diffusion transformers or vision-language-action (VLA) models Octo Model Team et al.
-[2024], Kim et al.
-[2024], could enhance representation learning, enabling better generalization, scalability, and transfer across tasks.
+Our simplification of the sensory input space increases training speed but may limit the model’s ability to generalize to visually complex or occlusion-prone construction scenarios, such as joint assembly in cluttered framing environments or multi-alignment tasks.
+Moreover, incorporating expressive state-of-the-art learning models, such as diffusion transformers or vision-language-action models (VLAs) Physical Intelligence et al.
+[2025], Gemini Robotics Team et al.
+[2025], enables better generalization, scalability, and transfer across tasks.
 These capabilities are particularly relevant for building-scale construction where robots must integrate multiple sensing modalities and operate across varied materials, such as combining timber, metal fasteners, and composite claddings in a single assembly sequence.
@@
 [2021].
-Future work could address these limitations by exploring hybrid learning paradigms that combine demonstration-based training with simulation pretraining, self-improvement cycles, or reinforcement learning, enabling policies to acquire physical reasoning skills and recover from unseen conditions TRI LBM Team et al.
-[2025].
-Additionally, emerging research on general-purpose robot foundation models, such as visuomotor foundation models and Large Behavior Models (LBMs) TRI LBM Team et al.
-[2025], Zhao et al.
-[2023], presents a promising direction for learning transferable policies across a broad range of dexterous manipulation tasks.
+Future work could address these limitations by exploring hybrid or hierarchical learning paradigms that combine demonstration-based training with simulation pretraining or reinforcement learning, enabling policies to acquire physical reasoning skills and recover from unseen conditions.
+Additionally, emerging research on general-purpose visuomotor foundation models such as VLAs or Large Behavior Models TRI LBM Team et al.
+[2025], Physical Intelligence et al.
+[2025] presents a promising direction for learning transferable policies across a broad range of dexterous manipulation tasks.
 Acknowledgments